### v1

# Decomposable Submodular Maximization in Federated Setting

Abstract
Submodular functions, as well as the sub-class of decomposable submodular functions, and their optimization appear in a wide range of applications in machine learning, recommendation systems, and welfare maximization. However, optimization of decomposable submodular functions with millions of component functions is computationally prohibitive. Furthermore, the component functions may be private (they might represent user preference function, for example) and cannot be widely shared. To address these issues, we propose a federated optimization setting for decomposable submodular optimization. In this setting, clients have their own preference functions, and a weighted sum of these preferences needs to be maximized. We implement the popular continuous greedy algorithm in this setting where clients take parallel small local steps towards the local solution and then the local changes are aggregated at a central server. To address the large number of clients, the aggregation is performed only on a subsampled set. Further, the aggregation is performed only intermittently between stretches of parallel local steps, which reduces communication cost significantly. We show that our federated algorithm is guaranteed to provide a good approximate solution, even in the presence of above cost-cutting measures. Finally, we show how the federated setting can be incorporated in solving fundamental discrete submodular optimization problems such as Maximum Coverage and Facility Location.
Federated Optimization, Constrained Optimization, Submodular Function
1 Introduction
Submodularity of a set function implies a natural diminishing returns property where the marginal benefit of any given element decreases as we select more and more elements. Formally, a set function $F\colon 2^{E}\to\mathbb{R}$ is submodular if for any $S\subseteq T\subseteq E$ and $e\in E\setminus T$ it holds that $F(S\cup\{e\})-F(S)\geq F(T\cup\{e\})-F(T).$ Decomposable submodular functions is an important subclass of submodular functions which can be written as sums of several component submodular functions: $F(S)=\sum_{i=1}^{N}f_{i}(S)$, for all $S\subseteq E$, where each $f_{i}:2^{E}\to\mathbb{R}$ is a submodular function on the ground set $E$ with $|E|=n$.
Decomposable submodular functions include some of the most fundamental and well-studied submodular functions such as max coverage, graph cuts, welfare maximization etc., and have found numerous applications in machine learning (Dueck and Frey, 2007; Gomes and Krause, 2010; Mirzasoleiman, Badanidiyuru, and Karbasi, 2016a, b; Mirzasoleiman, 2017), economics (Dobzinski and Schapira, 2006; Feige, 2006; Feige and Vondrák, 2006; Papadimitriou et al., 2008; Vondrák, 2008), social network (Kempe et al., 2003), and data summarization and recommender systems (Dueck and Frey, 2007; Gomes and Krause, 2010; Tschiatschek et al., 2014; Lin and Bilmes, 2011).
As a consequence of these wide-prevailing applications, maximizing decomposable submodular functions subject to different constraints has attracted significant theoretical/practical interest. The main approach in these cases is the centralized and sequential greedy approach.
The need for scalable and efficient optimization methods, which do not require collecting raw data in a central server and ensure secure information collection, is widespread in applications handling sensitive data such as medical data, web search queries, salary data, and social networks. In many such cases, individuals and companies are reluctant to share their data and collecting their data in a central server is a violation of their privacy. Moreover, collecting and storing all data on a single server or cluster is computationally expensive and infeasible for large-scale datasets, particularly when working with high-dimensional data or complex models. Thus, there is a widespread demand for scalable optimization algorithms that are both decentralized and prioritize privacy. Below are some examples that motivate the focus of this paper.
Example 1.1 (Welfare Maximization).
The welfare maximization problem aims to maximize the overall utility or welfare of a group of individuals or agents $a_{1},\dots,a_{N}$. In this problem, there is a set of items or goods $E$, and each individual $a_{i}$ has a certain preference or utility expressed as a submodular function $f_{i}:2^{E}\to\mathbb{R}_{+}$ that assigns a value to each combination of items. The goal is to partition $E$ into disjoint subsets $S_{1},\dots,S_{N}$ in order to maximize the social welfare $\sum_{i=1}^{N}f_{i}(S_{i})$. In an important special case, called Combinatorial Public Projects (Gupta et al., 2010; Papadimitriou et al., 2008), the goal is to find a subset $S\subseteq E$ of size at most $k$ that maximizes $F(S)=\sum_{i=1}^{N}f_{i}(S)$. This problem can be applied in different fields, such as resource allocation, public goods provision, market design, and has been intensively studied (Khot et al., 2008; Lehmann et al., 2006; Mirrokni et al., 2008). An optimal approximation algorithm is known for this problem in value oracle model in which it is required to have access to the value of $f_{i}(S)$ for each agent and any $S\subseteq E$ (Călinescu et al., 2011). However, in scenarios where agents are hesitant to disclose their data to a central server and storing all data on a single server is computationally infeasible, the demand for a decentralized and private submodular maximization algorithm becomes imperative.
Example 1.2 (Feature Selection).
Enabling privacy-protected data sharing among clinical centers is crucial for global collaborations. In the context of feature selection over sensitive medical data, consider geographically dispersed hospitals, denoted as $H_{1},\dots,H_{N}$, with each hospital maintaining its own data that it is unwilling to share. The goal is to identify a small subset $S$ of features that effectively classifies the target variable across the entire dataset over all hospitals. The demand for a decentralized and private feature selection algorithm is paramount to discover hidden patterns while preserving data ownership. By adopting a decentralized approach, hospitals can balance collaborative knowledge discovery and data privacy. One approach is to maximize a submodular function capturing the mutual information between features and the class labels (Krause and Guestrin, 2005).
Federated setting for learning and optimization. Federated setting (Konečnỳ et al., 2016; McMahan et al., 2017) is a novel and practical framework that addresses issues regarding privacy, data sharing, and centralized computation. On one hand, it is a distributed and collaborative approach that allows multiple parties, such as different organizations or devices, to train a shared model or collaboratively optimize an objective function while keeping their data locally. This approach helps to protect the privacy of data by ensuring that the raw data is never shared or moved outside of the individuals’ systems. Instead, only the model updates are exchanged and aggregated to improve the shared model and improve the objective value.
On the other hand, the federated framework reduces the amount of data that needs to be transferred and processed at any one time, which can significantly reduce the computational complexity of the overall process. Additionally, federated setting can also take advantage of the computational resources available at each party, such as the processing power of mobile devices or edge devices, which can further reduce computational load on the server. This way, federated learning can train models more efficiently, even with large-scale datasets and complex models, and provide a scalable solution for distributed learning.
Problem definition and setting. In this paper we introduce the problem of maximizing a submodular function in the federated setting. Let $E$ be a ground set of size $n$ and $c_{1},\dots,c_{N}$ be $N$ clients each of whom has a private interest over $E$. Each client’s interest is expressed as a submodular function. Let $f_{i}:2^{E}\to\mathbb{R}_{+}$ be the associated submodular function of the $i$-th client. A central server wants to solve the following constrained distributed optimization model
$\max_{S\in\mathcal{I}}\left\{F(S)=\sum_{i=1}^{N}p_{i}f_{i}(S)\right\},$ (1)
where $\mathcal{I}$ is the independent sets of a matroid $\mathcal{M}$ with ground set $E$, and $p_{i}$ are pre-defined weights such that $\sum\limits_{i=1}^{N}p_{i}=1$. For instance, they can be set to $1/N$, or the fraction of data owned by each client. The constraint implies sets of particular properties, e.g., subsets of size at most $k$. Note that, the unconstrained optimization is a special case of this.
In the optimization problem (1) the data can be massively distributed over the number of clients $N$, which can be huge. Moreover unlike the traditional distributed setting, in the federated setting the server does not have control over clients’ devices nor on how data is distributed. For example, when a mobile phone is turned off or WiFi access is unavailable, the central server will lose connection to this device. Furthermore, client’s objective can be very different depending on their local datasets. To minimize communication overhead and server computation load, the number of communication rounds from clients to servers need to be minimal.
Constraints. We formally discuss factors of efficiency and restrictions that should be considered.
  • 1.
    Privacy: One of the main appeals of decentralized and federated setting is to preserve the privacy. There are several models of privacy and security that have been considered in the literature such as Differential Privacy (DP) and Secure Aggregator (SecAgg), and a mix of these two. Single-server SecAgg is a cryptographic secure multi-party computation (MPC) that enables clients to submit vector inputs, such that the server (an aggregator) can only decipher the combined update, not individual updates. This is usually achieved via additive masking over a finite group (Bell et al., 2020; Bonawitz et al., 2016). Note that secure aggregation alone does not provide any privacy guarantees. To achieve a DP-type guarantee, noise can be added locally, with the server aggregating the perturbed local information via SecAgg. This user-level DP framework has recently been adopted in private federated learning (Agarwal et al., 2018, 2021; Kairouz et al., 2021; Chen et al., 2022a).
    In this paper we use Single-server SecAgg model of privacy, a dominant and well-established approach in the field. We leave other notions of privacy, such as a mix of DP and SecAgg, for future works. There has been a recent and concurrent progress towards this direction in terms of cardinality constraints (Wang et al., 2023).
  • 2.
    Communication and bit complexity: There are a few aspects to this, firstly the number of communication rounds between the server and clients should be as small as possible. Second, the information communicated between should require low bandwidth and they better require small bit complexity to encode.
  • 3.
    Convergence and utility: While the above impose strong restrictions, a good decentralized submodular maximization algorithm should not scarify the convergence rate by too much and should yield to an accurate and acceptable result in comparison to the centeralized methods.
Our contributions.
We present the first federated (constrained) submodular maximization algorithm that converges close to optimum guarantees known in centralized settings.
$\bullet$ We propose a decentralized version of the popular Continuous Greedy algorithm Federated Continuous Greedy (FedCG) and prove its convergence whenever the client functions are nonnegative monotone submodular achieving the optimal multiplicative approximation factor $(1-\nicefrac{{1}}{{e}})$ with a small additive.
$\bullet$ We incorporate important and practical scenarios that are relevant for federated setting such as partial client selection, low communication rounds and computation cost. We give rigorous theoretical guarantees under each scenarios matching the optimal multiplicative approximation of $(1-\nicefrac{{1}}{{e}})$ and small additive error.
$\bullet$ We introduce a new algorithm that serves as a discrete federated optimization algorithm for submodular maximization. Its convergence and applications to discrete problems such as Facility Location and Maximum Coverage are explored (Section 5).
1.1 Related work
FedAvg and its convergence.
The concept of Federated Learning (FL), introduced by McMahan et al. (2017), has found application in various domains such as natural language processing, computer vision, and healthcare. The popular FL algorithm, Federated Averaging (FedAvg), is an extension of Local SGD that aims to reduce communication costs in distributed settings (Gorbunov et al., 2021; Stich, 2019; Wang and Joshi, 2021; Yu et al., 2019). However, despite its practical benefits in addressing privacy, data heterogeneity, and computational constraints, it may not converge to a “good enough” solution in general (Pathak and Wainwright, 2020; Zhang et al., 2020), and analyzing the convergence of FedAvg and providing theoretical guarantees is challenging and necessitates making certain assumptions. Assumptions related to bounded gradients, convexity, Lipschitzness, statistical heterogeneity, and bounded variance of stochastic gradients for each client have been explored in recent works (Karimireddy et al., 2020; Li et al., 2020; Stich, 2019; Woodworth et al., 2020; Wang et al., 2019; Yu et al., 2019).
Decentralized / distributed submodular maximization. The main approach to submodular maximization is the greedy approach which in fact, in the centralized setting yields the tight approximation guarantee in various scenarios and constraints e.g., see (Nemhauser et al., 1978; Vondrák, 2008; Călinescu et al., 2011). Centeralized submodular maximization under privacy constraints is an active research area (Mitrovic et al., 2017; Rafiey and Yoshida, 2020; Chaturvedi et al., 2021). However the sequential nature of the greedy approach makes it challenging to scale it to massive datasets. This issue is partially addressed by the means of Map-Reduce style algorithms (Kumar et al., 2015) as well as several elegant algorithms in the distributed setting (Mirzasoleiman et al., 2016b; Barbosa et al., 2015). Recent work of Mokhtari et al. (2018b) ventures towards decentralized submodular maximization for continuous submodular functions. Note that in general continuous submodular functions are not convex nor concave and there has been a line of work to optimize continuous submodular functions using SGD methods (Hassani et al., 2017). Mokhtari et al. under several assumptions, such as assuming clients’ local objective functions are monotone, DR-submodular, Lipschitz continuous, and have bounded gradient norms, prove that Decenteralized Continuous Greedy algorithm yields a feasible solution with quality $O(1-1/\mathrm{e})$ times the optimal solution. The setting in (Mokhtari et al., 2018b) is fundamentally different from the federated setting in a sense that they require sharing gradient information of the clients with the server or with the neighboring nodes in an underlying graph. Perhaps the most closely related method to our work is due to Dadras et al. (2022); Zhang et al. (2022). Dadras et al. (2022) consider Frank-Wolfe Algorithm (Frank and Wolfe, 1956) in the federated setting and propose Federated Frank-Wolfe (FedFW) algorithm and analyze its convergence for both convex and non-convex functions and under the $L$-Lipschitzness and bounded gradients assumptions.
2 Preliminaries
Let $E$ denote the ground set and $|E|=n$. For a vector $\mathbf{x}\in\mathbb{R}^{|E|}$ and a set $S\subseteq E$, we use $\mathbf{x}(S)$ to denote $\sum_{e\in S}\mathbf{x}(e)$. We say a submodular function $f:2^{E}\to\mathbb{R}$ is monotone if $f(S)\leq f(T)$ for every $S\subseteq T\subseteq E$. Throughout this paper we assume $f(\emptyset)=0$.
Multilinear extension. The multilinear extension $\widehat{f}:[0,1]^{|E|}\to\mathbb{R}$ of a set function $f:\{0,1\}^{|E|}\to\mathbb{R}$ is
$\widehat{f}(\mathbf{x})$ $=\sum_{S\subseteq E}f(S)\prod\limits_{e\in S}\mathbf{x}(e)\prod\limits_{e\not\in S}(1-\mathbf{x}(e))=\mathbb{E}_{R\sim\mathbf{x}}[f(R)]$
where $R$ is the random subset of $E$ in which each element $e\in E$ is included independently with probability $\mathbf{x}(e)$ and is not included with probability $1-\mathbf{x}(e)$. We write $R\sim\mathbf{x}$ to denote that $R\subseteq E$ is a random set sampled according to $\mathbf{x}$.
Observe that for all $S\subseteq E$ we have $\widehat{f}(\mathbf{1}_{S})=f(S)$. For monotone non-decreasing submodular function $f$, $\widehat{f}$ has the following properties (Călinescu et al., 2011) that are crucial in analyses of our algorithms
  • 1.
    $\widehat{f}$ is monotone, meaning $\frac{\partial\widehat{f}}{\partial\mathbf{x}(e)}\geq 0$. Hence, $\nabla\widehat{f}(\mathbf{x})=(\frac{\partial\widehat{f}}{\partial\mathbf{x}(1)},\dots,\frac{\partial\widehat{f}}{\partial\mathbf{x}(n)})$ is a nonnegative vector.
  • 2.
    $\widehat{f}$ is concave along any direction $\mathbf{d}\geq\mathbf{0}$.
Note that $\frac{\partial\widehat{f}}{\partial\mathbf{x}(e)}=\mathbb{E}_{R\sim\mathbf{x}}[f(R\cup\{e\})-f(R\setminus\{e\})]$. That is the expected marginal contribution for $e$ where the expectation is taken over $R\subseteq E\setminus\{e\}$ sampled according to $\mathbf{x}$. By submodularity, for any $\mathbf{x}\in[0,1]^{n}$,
$|\nabla\widehat{f}(\mathbf{x})|_{\infty}\leq\max_{e\in E}f(\{e\}):=m_{f}.$ (2)
Matroids and matroid polytopes.
A pair $\mathcal{M}=(E,\mathcal{I})$ of a set $E$ and $\mathcal{I}\subseteq 2^{E}$ is called a matroid if 1) $\emptyset\in\mathcal{I}$, 2) $A\in\mathcal{I}$ for any $A\subseteq B\in\mathcal{I}$, and 3) for any $A,B\in\mathcal{I}$ with $|A|<|B|$, there exists $e\in B\setminus A$ such that $A\cup\{e\}\in\mathcal{I}$. We call a set in $\mathcal{I}$ an independent set. We sometimes abuse notation and use $S\in\mathcal{M}$. The rank function $r_{\mathcal{M}}\colon 2^{E}\to\mathbb{Z}_{+}$ of $\mathcal{M}$ is $r_{\mathcal{M}}(S)=\max\{|I|:I\subseteq S,I\in\mathcal{I}\}$. An independent set $S\in\mathcal{I}$ is called a base if $r_{\mathcal{M}}(S)=r_{\mathcal{M}}(E)$. We denote the rank of $\mathcal{M}$ by $r(\mathcal{M})$. The matroid polytope $\mathcal{P}(\mathcal{M})\subseteq\mathbb{R}^{E}$ of $\mathcal{M}$ is $\mathcal{P}(\mathcal{M})=\mathrm{conv}\{\mathbf{1}_{I}:I\in\mathcal{I}\}$ where $\mathrm{conv}$ denotes the convex hull. Or equivalently (Edmonds, 2001), $\mathcal{P}(\mathcal{M})=\left\{\mathbf{x}\geq\mathbf{0}:\mathbf{x}(S)\leq r_{\mathcal{M}}(S);~{}\forall S\subseteq E\right\}.$ The base polytope of $\mathcal{M}$ is given by $B(\mathcal{M})=\{\mathbf{y}\in\mathcal{P}(\mathcal{M})\mid\mathbf{y}(E)=r_{\mathcal{M}}(E)\}$. The extreme points of $B(\mathcal{M})$ are the characteristic vectors of the bases of $\mathcal{M}$.
The Continuous Greedy Algorithm.
Our algorithms for maximizing a submodular function in federated settings are based on the Continuous Greedy (CG) algorithm. Here we briefly explain this algorithm. The results mentioned are from (Călinescu et al., 2011; Vondrák, 2008). Let $\mathcal{M}=(E,\mathcal{I})$ be a matroid and $\mathcal{P}(\mathcal{M})$ be its matroid polytope of rank $r$, let $f$ be a nonnegative and monotone submodular function and $\widehat{f}$ be its multilinear extension. CG starts with $\mathbf{x}^{(0)}=\mathbf{0}$. For every $t\in\{0,1,2,\dots,T-1\}$ it computes $\mathbf{x}^{(t+1)}$ using the following update step $\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\mathbf{v}^{(t)}$, where $\mathbf{v}^{(t)}=\operatornamewithlimits{argmax}_{\mathbf{w}\in\mathcal{P}}\langle\mathbf{w},\nabla\widehat{f}(\mathbf{x}^{(t)})\rangle$. For $\texttt{OPT}=\max_{\mathbf{x}\in\mathcal{P}}\widehat{f}(\mathbf{x})$ we have $(1-(1-\eta)^{T})\texttt{OPT}\leq\widehat{f}(\mathbf{x}^{(T)})+C\eta^{2}/2$. Here, the constant $C$ depends on the Lipschitz of the function, and $\mathbf{x}^{(T)}\in\mathcal{P}$ as it is a convex combination of vectors from the polytope. Moreover, for $\eta=1/T$ and large enough $T$ we get $(1-1/\mathrm{e})\texttt{OPT}\leq\widehat{f}(\mathbf{x}^{(T)})+\epsilon$.
Given $\mathbf{x}^{(T)}\in\mathcal{P}$, there are rounding procedures to obtain $S\in\mathcal{I}$ such that $\widehat{f}(\mathbf{x}^{(T)})\leq f(S)$. The approximation factor $1-1/\mathrm{e}$ is the best possible assuming $\mathbf{P}\neq\mathbf{NP}$ (Feige, 1998).
3 Federated Continuous Greedy
In this section we propose our Federated Continuous Greedy (FedCG) method. We start with a simplistic scenario of federated model with full participation which already shows some of the challenges that we have to overcome before delving into the partial participation model which is more computationally feasible.
Consider optimization problem (1) where each $f_{i}$ is a nonnegative monotone submodular function, $\sum_{i=1}^{N}p_{i}=1$, and $\mathcal{I}$ is the independent sets of matroid $\mathcal{M}=(E,\mathcal{I})$ of rank $r$.
Bit complexity and accuracy trade-off.
For every client $i$ and every $\mathbf{x}\in\mathcal{P}$, the vector $\mathbf{v}_{i}=\operatornamewithlimits{argmax}_{\mathbf{w}\in\mathcal{P}}\langle\mathbf{w},\nabla\widehat{f_{i}}(\mathbf{x})\rangle$ is determined by maximizing a linear function $\langle\mathbf{w},\nabla\widehat{f_{i}}(\mathbf{x})\rangle$ over $\mathcal{P}$. In the case of a matroid polytope $\mathcal{P}$, this problem can be solved very efficiently. We can assume that $\mathbf{v}_{i}$ is a vertex of $\mathcal{P}$ and furthermore, since $\nabla\widehat{f_{i}}$ is a nonnegative vector, that this vertex corresponds to a base of matroid $\mathcal{M}$. Hence, without loss of generality $\mathbf{v}_{i}$ is the indicator vector of a base with $r$ ones and $n-r$ zeros. Thus it can be encoded using $O(r\log(n))$ bits, sublinear in the size of the ground set. On the other hand, the vector $\nabla\widehat{f_{i}}(\mathbf{x})$ itself requires $\tilde{O}(n)$ bits for encoding. In what follows we will see how restricting the bit complexity effects the accuracy, and the amount of computation that server should do.
FedCG with full participation.
First consider the case where clients can send their gradients. We proceeds in rounds. Initially $\mathbf{x}^{(0)}=0$. On the $t$-th round, first, the central server broadcasts the latest model $\mathbf{x}^{(t)}$ to all clients. Each client $i$ after receiving the update sets $\mathbf{x}_{i}^{(t)}=\mathbf{x}^{(t)}$ and computes $\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})$. The server then aggregates local information via SecAgg, and computes $\nabla\widehat{F}(\mathbf{x}^{(t)})=\sum_{i=1}^{N}p_{i}\nabla\widehat{f}_{i}({\mathbf{x}^{(t)}})$. After receiving $\nabla\widehat{F}(\mathbf{x}^{(t)})$, the server computes $\mathbf{v}^{(t)}=\operatornamewithlimits{argmax}_{\mathbf{w}\in\mathcal{P}}\langle\mathbf{w},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ by maximizing a linear function subject to the matroid constraint and produces the new global model with learning rate $\eta$:
$\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\mathbf{v}^{(t)}$ (3)
It is clear that, similar to the centeralized CG, for large enough $T$ we get $F(\mathbf{x}^{(T)})\geq(1-\nicefrac{{1}}{{\mathrm{e}}})\texttt{OPT}$. Note that this simple framework has an advantage over centralized CG, it is taking advantage of the computational resources available at each client.
Second consider the more challenging case where clients can send at most $\tilde{O}(r)$ bits information. We see how this restriction effects the accuracy. Our algorithm, FedCG, proceeds in rounds. Initially $\mathbf{x}^{(0)}=0$. On the $t$-th round, first, the central server broadcasts the latest model $\mathbf{x}^{(t)}$ to all clients. Each client $i$ after receiving the update sets $\mathbf{x}_{i}^{(t)}=\mathbf{x}^{(t)}$ and performs one step of continuous greedy approach to find a direction that best aligns with her local gradient:
$\mathbf{v}_{i}^{(t)}\leftarrow\operatornamewithlimits{argmax}_{\mathbf{v}\in\mathcal{P}}\langle\mathbf{v},\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t)})\rangle$ (4)
Lastly, clients send their update directions $\mathbf{v}_{1}^{(t)},\dots,\mathbf{v}_{N}^{(t)}$ to the secure aggregator to compute $\Delta^{(t)}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)}$. After receiving $\Delta^{(t)}$, the server produces the new global model with learning rate $\eta$:
$\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\Delta^{(t)}$ (5)
Even in this unrealistic setting where all clients participate in each round the convergence analysis requires new insights. In order to provide an approximation guarantee for our algorithm we shall obtain a lower bound on the function value improvement by taking the direction $\Delta^{(t)}$, that is providing a lower bound for $\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$. However, each $\mathbf{v}_{i}^{(t)}$ is the projection of the local gradient into the matroid polytope and it does not carry information about the magnitude of the expected marginal contributions i.e. $\|\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\|$. Without assuming an assumption on the heterogeneity of local functions one can construct examples where a single element and corresponding client’s marginal contribution are significantly more dominant than others and hence taking direction $\Delta^{(t)}$ results in a very bad approximation guarantee.
We therefore need an assumption that acts as a tool in constraining the level of heterogeneity which poses a significant obstacle in federated optimization. A common way to handle heterogeneity is to impose a bound on the magnitude of gradients over each clients’ local function i.e $\|\nabla\widehat{f}_{i}(\mathbf{x})\|\leq\gamma$. This types of assumption is not only common in the literature regarding submodular function maximization in decentralized settings (Mokhtari et al., 2018b; Zhang et al., 2020), but also in studies of convex and non-convex optimization in federated learning (Chen et al., 2022b; Dadras et al., 2022; Karimireddy et al., 2020; Yu et al., 2019; Li et al., 2020).
In the case of submodular functions, each coordinate of the gradient of the multilinear extension corresponds to the marginal gain of adding one single element. In this paper we impose the following assumption which is much more relaxed than assuming a bound on the magnitude of the gradients from each client.
Assumption 3.1.
For all $i=1,\dots,N$ and $t=1,\dots,T$ we have $|\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$. Note that, by submodularity and monotonicity we have
$\max_{t\in[T]}\gamma_{t}\leq 2\max_{i}\max_{e\in E}f_{i}(\{e\})=2\max_{i}m_{f_{i}}.$ (6)
Note that monotonicity of $f_{i}$ implies that for every $\mathbf{x}\leq\mathbf{y}$ corrdinate-wise, it holds that $\widehat{f}_{i}(\mathbf{x})\leq\widehat{f}_{i}(\mathbf{x})$. Additionally, gradients are antitone i.e., for every $\mathbf{x}\leq\mathbf{y}$ corrdinate-wise, it holds that $\nabla\widehat{f}_{i}(\mathbf{x})\geq\nabla\widehat{f}_{i}(\mathbf{y})$. Thus as the algorithm advances and $t$ grows, $\gamma_{t}$’s change but never exceed the upper bound in (6). Additionally, in numerous instances, $\max_{t\in[T]}\gamma_{t}$ is relatively small. For instance, in Max Coverage problem each $f_{i}(S)$ is either 0 or 1 , depending if a client is covered by $S$ or not, thus for this problem $\gamma=1$.
We now have enough ingredients to prove the following convergence theorem for the case where all clients participate in every communication round. Let $D=\sum_{t=1}^{T}\gamma_{t}$
Theorem 3.2 (Full participation).
Let $\mathcal{M}$ be a matroid of rank $r$ and $\mathcal{P}$ be its matroid polytope. Under the full participation assumption and Assumption 3.1, for every $\eta>0$, Algorithm 1 returns a $\mathbf{x}^{(T)}\in\mathcal{P}$ such that
$\left(1-(1-\eta)^{T}\right)\texttt{OPT}\leq\widehat{F}(\mathbf{x}^{(T)})+\eta r\sum_{t=1}^{T}\gamma_{t}+\frac{T\eta^{2}r^{2}m_{F}}{2}$
In particular, for large enough $T$, setting $\eta=1/T$, Algorithm 1 requires at most $\tilde{O}(r)$ bits of communication per user per round ($\tilde{O}(NTr)$ in total) and obtains
$(1-1/\mathrm{e})\texttt{OPT}\leq\widehat{F}(\mathbf{x}^{(T)})+\left(\frac{rD}{T}+\frac{r^{2}m_{F}}{2T}\right).$
Algorithm 1 Federated Continuous Greedy (FedCG)
1:  Input: Matroid polytope $\mathcal{P}$, number of communication rounds $T$, learning rate $\eta$, and $K$.
2:  $\mathbf{x}^{(0)}=\mathbf{0}$
3:  for $t=0$ to $T-1$ do
4:     Server selects a subset of $K$ active clients $A^{(t)}$ according to Client Sampling Scheme, and sends $\mathbf{x}^{(t)}$ to them.
5:     for Client $i$ in $A^{(t)}$ in parallel  do
6:        $\mathbf{v}_{i}^{(t)}\leftarrow\operatornamewithlimits{argmax}_{\mathbf{v}\in\mathcal{P}}\langle\mathbf{v},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$
7:        Send $\mathbf{v}_{i}^{(t)}$ back to the secure aggregator.
8:     end for
9:     SecAgg: $\Delta^{(t)}=\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\mathbf{v}_{i}^{(t)}$
10:     Server updates: $\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\Delta^{(t)}$
11:  end for
12:  Apply a proper rounding scheme on $\mathbf{x}^{(T)}$ to obtain a solution for (1)
3.1 Partial participation and client selection
Client sampling in the FL optimization framework is imperative for various practical reasons, including the following:
$\bullet$ Large scale and dynamic nature. In real-world applications, a server usually serves several billions of devices/clients who can join or leave the federated optimization system due to several reasons like intermittent connectivity, technical issue, or simply based on their availability or preferences. Hence, it is computationally inefficient and often impossible to get updates from all clients.
$\bullet$ Communication and bandwidth. On one hand, waiting for the slowest client to finish can increase the expected round duration as the number of participating clients per training increases, a phenomenon known as “straggler’s effect”. On the other hand, communication can be a primary bottleneck for federated settings because of clients bandwidth limitation and the possibility of server throttling.
$\bullet$ Small models and redundancy. It is often the case that FL models are small because of clients limited computational power or memory, it therefore is unnecessary to train an FL model on billions of clients. Note that for optimization problem (1) in many practical scenarios we are dealing with small models in comparison to the number of clients, this is because of the matroid constraint or simply because the size of the ground set is much smaller than the number of clients.
Here we discusse our sampling scheme which crucially does not violate clients’ privacy. The overall algorithm is given in Algorithm 1.
Unbiased Client Sampling Scheme.
At each communication round the server chooses an active client from $i\in[N]$ with probability $p_{i}$, and repeats this process $K$ times to obtain a multiset $A^{(t)}$ of size $K$ which may contain a client more than once. Then the aggregation step is $\Delta^{(t)}=\frac{1}{K}\sum\limits_{i\in A^{(t)}}\mathbf{v}_{i}^{(t)}$ where $\mathbf{v}_{i}^{(t)}$ defined in (4).
The next lemma shows that this sampling scheme is unbiased and in expectation the average update from chosen clients $A^{(t)}$ is equal to the average update from all clients.
Lemma 3.3 (Unbiased sampling scheme).
For Client Sampling Scheme, we have $\mathbb{E}_{A^{(t)}}\left[\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$.
The following lemma allows us to bound the variance which in turn helps to provide our convergence guarantees. To show the following result, it is required to upper bound the difference between the improvement on the function value by taking the direction suggested by a selected client versus taking the direction obtained by averaging all the directions from the clients. That is bounding $|\langle\mathbf{v}_{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\sum_{i=1}^{N}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|$, for selected client $s$ in $A^{(t)}$. This in turn needs providing an upper bound for $\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{v}_{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$, for all $i\neq s$. At the heart, our proof relies on the properties of multilinear extensions of local submodular functions, the fact that each $\mathbf{v}_{i}^{(t)}$ corresponds to a base of the matroid, and Assumption 3.1.
Lemma 3.4 (Bounded variance).
Using Client Sampling Scheme we have $\mathrm{Var}(\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle)\leq 36r^{2}\gamma_{t}^{2}/K$.
Armed with the above lemmas and concentration inequalities e.g., Chebyshev’s inequality, we can prove the convergence of Algorithm 1. This essentially is done by bounding the error introduced by the decentralized setting and carefully carrying the error through the analysis.
Theorem 3.5.
Let $\mathcal{M}$ be a matroid of rank $r$ and $\mathcal{P}$ be its matroid polytope. Using Client Sampling Scheme, for every $\eta,\delta>0$, Algorithm 1 returns a $\mathbf{x}^{(T)}\in\mathcal{P}$ so that with probability at least $1-\delta$
$\left(1-(1-\eta)^{T}\right)\texttt{OPT}\leq$
$~{}~{}~{}~{}~{}\widehat{F}(\mathbf{x}^{(T)})+\eta\left(r\sum_{t=1}^{T}\gamma_{t}+\frac{6r\sum_{t=1}^{T}\gamma_{t}}{\sqrt{K\delta/T}}\right)+\frac{T\eta^{2}r^{2}m_{F}}{2}$
In particular, by setting $\eta=1/T$, Algorithm 1 requires at most $\tilde{O}(r)$ bits of communication per user per round ($\tilde{O}(KTr)$ in total) and yields
$\left(1-1/\mathrm{e}\right)\texttt{OPT}\leq\widehat{F}(\mathbf{x}^{(T)})+\left(\frac{rD}{T}+\frac{6rD}{\sqrt{TK\delta}}+\frac{r^{2}m_{F}}{2T}\right)$
4 Practical Federated Continuous Greedy
One of the main considerations in federated optimization is the number of communication rounds. In this section, we show how Algorithm 1 can be further improved to reduce communication rounds while simultaneously incorporating partial participation.
Algorithm 2 Practical FedCG (FedCG+)
1:  Input: Matroid polytope $\mathcal{P}$, number of communication rounds $T/\tau$, server’s learning rate $\eta$, $\sigma,\delta>0$, and $K$.
2:  $\mathbf{x}^{(0)}=\mathbf{0}$, $m=O(\log{(TK/\delta)}/\sigma^{2})$
3:  for $t=0,\tau,2\tau,\dots,(T-1)/\tau$ do
4:     Server selects a subset of $K$ active clients $A^{(t)}$ according to Client Sampling Scheme, and sends $\mathbf{x}^{(t)}$ to them.
5:     for Client $i$ in $A^{(t)}$ in parallel  do
6:        $\mathbf{x}_{i}^{(t,0)}\leftarrow\mathbf{x}^{(t)}$
7:        for $j=0,\dots,\tau-1$ do
8:           Randomly sample $m$ sets $\zeta_{i}^{(t,j)}=\{R_{i}^{(t,j,1)},\dots,R_{i}^{(t,j,m)}\}$ according to $\mathbf{x}_{i}^{(t,j)}$
9:           Let $\nabla\widetilde{f}_{i}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})$ be the estimate of $\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t,j)})$
10:           $\widetilde{\mathbf{v}}_{i}^{(t,j)}\leftarrow\operatornamewithlimits{argmax}_{\mathbf{v}\in\mathcal{P}}\langle\mathbf{v},\nabla\widetilde{f}_{i}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})\rangle$
11:           $\mathbf{x}_{i}^{(t,j+1)}\leftarrow\mathbf{x}_{i}^{(t,j)}+\widetilde{\mathbf{v}}_{i}^{(t,j)}/\tau$
12:        end for
13:        $\widetilde{\Delta}_{i}^{(t+\tau)}\leftarrow\mathbf{x}_{i}^{(t,\tau)}-\mathbf{x}_{i}^{(t,0)}$Local model change}
14:        Send $\widetilde{\Delta}_{i}^{(t+\tau)}$ back to the secure aggregator.
15:     end for
16:     SecAgg: $\widetilde{\Delta}^{(t+\tau)}=\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\widetilde{\Delta}_{i}^{(t+\tau)}$.
17:     Server updates: $\mathbf{x}^{(t+\tau)}\leftarrow\mathbf{x}^{(t)}+\eta\widetilde{\Delta}^{(t+\tau)}$
18:  end for
19:  Apply a proper rounding scheme on $\mathbf{x}^{(T)}$ to obtain a solution for (1)
Algorithm description. Initially $\mathbf{x}^{(0)}=0$. On the $t$-th round of the Practical Federated Continuous Greedy (FedCG+), the central server first broadcasts the latest model $\mathbf{x}^{(t)}$ to a subset of active clients of size $K$ denoted by $A^{(t)}$. Next, each client $i\in A^{(t)}$ sets $\mathbf{x}_{i}^{(t,0)}=\mathbf{x}^{(t)}$ and performs $\tau$ steps of continuous greedy approach locally. More precisely, let $j\in\{0,1,\dots,\tau-1\}$ and $\mathbf{x}_{i}^{(t,j)}$ denote the $i$-th client’s local model at communication round $t$ and local update step $j$, then the local updates are
$\widetilde{\mathbf{v}}_{i}^{(t,j)}\leftarrow\operatornamewithlimits{argmax}_{\mathbf{v}\in\mathcal{P}}\langle\mathbf{v},\nabla\widetilde{f}_{i}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})\rangle;$ (7)
$\mathbf{x}_{i}^{(t,j+1)}=\mathbf{x}_{i}^{(t,j)}+\widetilde{\mathbf{v}}_{i}^{(t,j)}/\tau$ (8)
Here $\zeta_{i}^{(t,j)}$ is a set of subsets from the ground set $E$ sampled according to $\mathbf{x}_{i}^{(t,j)}$, and $\nabla\widetilde{f}_{i}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})\in\mathbb{R}_{\geq 0}^{n}$ is an estimation of the gradient $\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t,j)})$ (more on this later).
After $\tau$ steps of local update, the $i$-th client from $A^{(t)}$ send her update $\widetilde{\Delta}_{i}^{(t+\tau)}=\mathbf{x}_{i}^{(t,\tau)}-\mathbf{x}_{i}^{(t,0)}$ to the secure aggregator to compute $\widetilde{\Delta}^{(t+\tau)}=\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\widetilde{\Delta}_{i}^{(t+\tau)}$. Note that each $\widetilde{\Delta}_{i}^{(t+\tau)}$ belongs to $\mathcal{P}$ since it is a convex combination of vectors $\widetilde{\mathbf{v}}_{i}^{(t,j)}\in\mathcal{P}$. However, $\widetilde{\Delta}_{i}^{(t+\tau)}$ may not be an integral vector and in the worst case $\tilde{O}(n)$ bits are required to encode it. After receiving $\widetilde{\Delta}^{(t+\tau)}$, the server produces the new global model with learning rate $\eta$:
$\mathbf{x}^{(t+\tau)}\leftarrow\mathbf{x}^{(t)}+\eta\widetilde{\Delta}^{(t+\tau)}$ (9)
Gradient estimation. Evaluating the multilinear extension involves summing over all subsets $S$ of $E$, there are $2^{|E|}$ such subsets. However, recall $\partial\widehat{f}/\partial\mathbf{x}(e)=\mathbb{E}[f(R\cup\{e\})]-\mathbb{E}[f(R\setminus\{e\})]$ where $R\subseteq E$ is a random subset sampled according to $\mathbf{x}$. Hence a simple application of Chernoff’s bound tells us by sampling sufficiently many subsets we can obtain a good estimation of $\nabla\widehat{f}(\mathbf{x})$ (Călinescu et al., 2011; Vondrák, 2008) (more details in the Appendix). For large enough $m$, let $\zeta_{i}^{(t,j)}=\{R_{i}^{(t,j,1)},\dots,R_{i}^{(t,j,m)}\}$ be subsets of $E$ that are sampled independently according to $i$-th client’s local model $\mathbf{x}_{i}^{(t,j)}$. Let $\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})$ denote the stochastic approximation of $\nabla\widehat{f_{i}}(\mathbf{x}_{i}^{(t,j)})$. Then with probability $1-\delta$ we have $\|\nabla\widehat{f_{i}}(\mathbf{x}_{i}^{(t,j)})-\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})\|\leq\sigma$.
In convergence analyses of our algorithm there are two sources of randomness, one in the sampling schemes for client selection and the other in data sampling at each clients local data to estimate the gradients. In Algorithm 2 there are $\frac{T}{\tau}$ communications rounds between the server and clients. We define $\mathcal{I}_{\tau}=\{\tau i\mid i=1,2,3,\dots\}$ to denote the set of communication rounds with the server. Similar to Lemma 3.3:
Lemma 4.1 (Unbiased sampling scheme).
For Client Sampling Scheme, at every communication round $t+\tau\in\mathcal{I}_{\tau}$, we have $\mathbb{E}_{A^{(t)}}\left[\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\langle\sum_{i=1}^{N}p_{i}\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$.
Bounding the variance of $\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ is much more delicate than in Lemma 3.4. The are two main reasons, one is the deviation caused by local steps, second is the deviation from the true $\nabla\widehat{F}(\mathbf{x}^{(t)})$ caused by gradient estimations. To handle the deviation cause by local steps we assume all $\widehat{f}_{i}$ have $L$-Lipschitz continuous gradients i.e., $\|\nabla\widehat{f}_{i}(\mathbf{x})-\nabla\widehat{f}_{i}(\mathbf{y})\|\leq L\|\mathbf{x}-\mathbf{y}\|$, $\forall\mathbf{x},\mathbf{y}\in\mathcal{P}$. In fact as shown in Lemma 3 of (Mokhtari et al., 2018a), for each $\widehat{f}_{i}$ and local model $\mathbf{x}_{i}^{(t,j+1)}=\mathbf{x}_{i}^{(t,j)}+\widetilde{\mathbf{v}}_{i}^{(t,j)}/\tau$, it holds that
$\left\|\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t,j+1)})-\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t,j)})\right\|\leq\frac{m_{f_{i}}\sqrt{r}}{\tau}\left\|\widetilde{\mathbf{v}}_{i}^{(t,j)}\right\|\leq\frac{m_{f_{i}}r}{\tau}$
The factor $-m_{f_{i}}$ is in fact a lower bound on the entries of the Hessian matrix. That is $\frac{\partial\widehat{f}_{i}}{\partial\mathbf{x}(i)\partial\mathbf{x}(j)}\geq-m_{f_{i}}$ (Hassani et al., 2017).
Let $Q=\sum_{t\in I_{\tau}}L_{t}$ where $L_{t}$ is such that for all $i\in[N]$ and $j\in[\tau]$ it holds
$\left\|\nabla\widehat{f_{i}}(\mathbf{x}_{i}^{(t,j)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\right\|\leq L_{t}\|\mathbf{x}_{i}^{(t,j)}-\mathbf{x}^{(t)}\|$
Observe that $L_{t}$ is upper bounded by $\max_{i}m_{f_{i}}\sqrt{r}$.
Lemma 4.2 (Bounded variance).
For $t+\tau\in\mathcal{I}_{\tau}$ we have $\mathrm{Var}(\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle)\leq\frac{1}{K}\left(6r\gamma_{t}+2(\sigma r+L_{t}r^{1.5})\right)^{2}$.
While the variance can be made arbitrary small by sampling more clients, the additive error caused by local steps cannot be controlled by sampling more clients. This shows up in the next theorem.
Theorem 4.3.
Let $\mathcal{M}$ be a matroid of rank $r$ and $\mathcal{P}$ be its matroid polytope. Using Client Sampling Scheme, for every $\eta,\delta>0$, Algorithm 2 returns a $\mathbf{x}^{(T/\tau)}\in\mathcal{P}$ such that with probability at least $1-\delta$ it holds
[…]
In particular, for $\eta=\tau/T$, Algorithm 2 has at most $\tilde{O}(n)$ bits of communication per user per round ($\tilde{O}(KTn/\tau)$ in total) and yields
$\left(1-1/\mathrm{e}\right)\texttt{OPT}\leq\widehat{F}(\mathbf{x}^{(T)})+\frac{\tau r^{2}m_{F}}{2T}+$
$\frac{rD\tau}{T}+2\sigma r(1+\frac{\sqrt{\tau}}{\sqrt{KT\delta}})+\frac{2r^{1.5}Q\tau}{T}+\frac{\sqrt{\tau}(6Dr+2r^{1.5}Q)}{\sqrt{TK\delta}}$
5 Discrete problems in federated setting
Algorithm 3 Discrete Federated Submodular Maximization
1:  Input: Matroid $\mathcal{M}$ of rank $r$, importance factors $\{w_{i}\}_{i=1}^{N}$, $\varepsilon\in(0,1)$.
2:  $S\leftarrow\emptyset$, $\kappa\leftarrow\tilde{O}(rn/\varepsilon^{2})$
3:  for $t=0$ to $r-1$ do
4:     Server sends $S$ to all clients.
5:     for each client $i$ in parallel do
6:        $\kappa_{i}\leftarrow\min\{\kappa\cdot w_{i},1\}$
7:        for all $e$ such that $S\cup\{e\}\in\mathcal{M}$ do
8:           $\Delta_{i}[e]\leftarrow(f_{i}(S\cup\{e\})-f_{i}(S))/\kappa_{i}$
9:        end for{/* Randomized Response */}
10:        With probability $\kappa_{i}$ sends $\Delta_{i}$ to the secure aggregator
11:        With probability $1-\kappa_{i}$ does nothing
12:     end for
13:     SecAgg: $\Delta^{(t)}$ (the sum of $\Delta_{i}$ it has received in this round.)
14:     Server updates: $S\leftarrow S\cup\{\operatornamewithlimits{argmax}_{e\in E\setminus S}\Delta^{(t)}[e]\}$
15:  end for
16:  Output: $S$
While parallel SGD and continuous methods such as ours in this paper are commonly used as the main tool in federated optimization, we introduce a rather discrete approach to the field and believe it will find further applications. Our approach is inspired by recent works of Rafiey and Yoshida (2022); Kenneth and Krauthgamer (2023) on a seemingly unrelated topic. Rafiey and Yoshida (2022) introduced a method to sparsify a sum of submodular functions in a centralized setting which was improved by (Kenneth and Krauthgamer, 2023). We tailor their approach to the federated setting and discuss its effectiveness for discrete problems such as Facility Location and Maximum Coverage problems. Recall the optimization problem:
$\max_{S\in\mathcal{I}}\left\{F(S)=\sum_{i=1}^{N}f_{i}(S)\right\}$
where $\mathcal{M}=(E,\mathcal{I})$ is a matroid of rank $r$. At the heart of this approach is for clients to know their “importance” without sharing sensitive information. In the monotone case, the importance factor for client $i$ is defined as
$w_{i}=\max_{e\in E}\frac{f_{i}(\{e\})}{F(\{e\})}.$
In several cases such as Max Facility Location and Maximum Coverage computing the importance factor can be done efficiently and with constant number of communication rounds. For now, let us continue by assuming each client knows its own importance factor.
Algorithm 3 description. Let $\varepsilon\in(0,1)$ and set $S=\emptyset$. The server gradually adds elements to $S$ for only $r$ rounds. At round $t$ the central server broadcasts the current set $S$ to all clients (or a subset of active clients). Then each client $i$ computes the marginal contribution of each element from $E\setminus S$ to its local function $f_{i}$
$\Delta_{i}[e]=f_{i}(S\cup\{e\})-f_{i}(S);\quad\forall e:S\cup\{e\}\in\mathcal{I}$
Let $\kappa=\tilde{O}(rn/\varepsilon^{2})$. Then each client sends its update in a randomized response manner, that is, the $i$-th client with probability $\kappa_{i}=\min\{1,\kappa\cdot w_{i}\}$ sends the scaled vector $(\frac{1}{\kappa_{i}})\Delta_{i}$ to the secure aggregator, and with the complement probability does not send anything. The secure aggregator computes $\Delta^{(t)}$; the sum of the update vectors it has received. The server updates set $S$:
$S\leftarrow S\cup\{\operatornamewithlimits{argmax}_{e:S\cup\{e\}\in\mathcal{M}}\Delta^{(t)}[e]\}$ (10)
The following theorem follows from the results in (Rafiey and Yoshida, 2022; Kenneth and Krauthgamer, 2023).
Theorem 5.1.
For every $\varepsilon\in(0,1)$, Algorithm 3, with probability at least $1-1/n$, returns a subset $S\in\mathcal{I}$ such that $(1-\frac{1}{\mathrm{e}}-\varepsilon)\texttt{OPT}\leq F(S).$ Moreover, at each round $\tilde{O}(kn^{2}/\varepsilon^{2})$ clients participate.
5.1 Discrete examples
We consider two well-studied problems, namely Facility Location and Maximum Coverage problems and discuss the details of how to compute importance factors in federated setting efficiently and prove that the expected number of clients participating in each round is small for these two problems.
5.1.1 Facility Location Problem
Let $\mathcal{C}$ be a set of $N$ clients and $E$ be a set of facilities with $|E|=n$. For $c:\mathcal{C}\times E\to\mathbb{R}$ let the $i$-th client’s score function over a subset of facilities be $f_{i}(A)=\max_{j\in A}c(i,j)$. The objective for Max Facility Location is
$\max_{S\subseteq E,|S|\leq k}\left\{F(A)=\sum_{i=1}^{N}\max_{j\in A}c(i,j)\right\}$
For each client $i$ the importance factor is $w_{i}=\max\limits_{j\in E}\frac{c(i,j)}{F(\{j\})}$.
It is straightforward to see each client can compute its corresponding importance factor in a federated setting and using a secure aggregator without sharing its data with other clients. Each client $i$ sends a vector $(c(i,1),\dots,c(i,n))$ to the server and by simply summing up these vectors the server has a histogram over facilities. This histogram is then broadcasts to the clients where they can compute their own importance factor (see Algorithm 4). This requires only two communication rounds. Having $w_{i}$ on hand we can execute Algorithm 3 for Max Facility Location problem. Note that at each round of Algorithm 3 the expected number of clients participating is $\tilde{O}(kn^{2}/\varepsilon^{2})$, independent of $N$.
5.1.2 Maximum Coverage Problem
Let $\mathcal{C}=\{C_{1},\dots,C_{N}\}$ be a set of clients and $E=\{G_{1},\dots,G_{n}\}$ be a family of sets where each $G_{i}\subseteq\mathcal{C}$ is a group of clients. Given a positive integer $k$, in the Max Coverage problem the objective is to select at most $k$ groups of clients from $E$ such that the maximum number of clients are covered, i.e., the union of the selected groups has maximal size. One can formulate this problem as follows. For every $i\in[N]$ and $A\subseteq[n]$ define $f_{i}(A)$ as
$f_{i}(A)=\begin{cases}1&\text{if there exists $a\in A$ such that$C_{i}\in G_{a}$},\\ 0&\text{otherwise.}\end{cases}$
Note that $f_{i}$’s are monotone and submodular. Furthermore, define $F(A)=\sum_{i\in[N]}f_{i}(A)$ which is monotone and submodular as well. Now the Max Coverage problem is equivalent to
$\max_{A\subseteq[n],|A|\leq k}\left\{F(A)=\sum_{i\in[N]}f_{i}(A)\right\}$ (11)
For each client $C_{i}$, its importance factor $w_{i}$ is $\max\limits_{\begin{subarray}{c}G_{a}\in E,C_{i}\in G_{a}\end{subarray}}\frac{1}{|G_{a}|}$.
Similar to the case of Facility Location the importance factors can be computed in federated setting by having the membership histograms over the groups (see Algorithm 5). Applying Algorithm 3 yields $(1-1/e-\varepsilon)$ approximation guarantee where at each round the expected number of participating clients is $\tilde{O}(kn^{2}/\varepsilon^{2})$, independent of $N$.
6 Conclusions and Future Work
We present FedCG, the first algorithm for decomposable submodular maximization in federated setting under matroid constraints. FedCG is based on the continuous greedy algorithm and achieves the best possible approximation factor i.e., $1-1/\mathrm{e}$ under mild assumptions even faced with client selection and low communication rounds. Additionally, we introduce a new federated framework for discrete problems. Our work leads to many interesting directions for future work, such as providing stronger privacy guarantee using differentially private methods developed for submodular maximization (Mitrovic et al., 2017; Rafiey and Yoshida, 2020; Chaturvedi et al., 2021), finding a lower bound on the bit complexity, and improving the additive error of our algorithms.
Acknowledgement
I would like to thank Arya Mazumdar and Barna Saha for their many useful discussions and for reading several drafts of this paper. I would also like to thank Yuichi Yoshida for his valuable discussions and inspiring comments on the paper. Additionally, I am grateful to Nazanin Mehrasa, Wei-Ning Chen, Heng Zhu, and Quanquan C. Liu for insightful comments on this work.
Appendix A Preliminary results and proof of convergence for full participation
Quick overview.
At the heart of the analyses for the approximation ratio of the Centralized Continuous Greedy is to show that at each iteration the algorithm reduces the gap to the optimal solution by a significant amount [Călinescu et al., 2011]. We follow the same general idea although there are several subtleties that we should address. Mainly, it is required to compare the improvement we obtain by taking direction $\mathbf{v}$ versus the improvement obtained by taking direction $\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}$ where $\mathbf{v}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{F}(\mathbf{x})\rangle$ and $\mathbf{v}_{i}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{f}_{i}(\mathbf{x})\rangle$. To establish a comparison between the two, we need several intermediate lemmas.
We start off by focusing on providing an upper bound on $\texttt{OPT}-\widehat{F}(\mathbf{x})$.
Lemma A.1.
Let $F=\sum_{i=1}^{N}p_{i}f_{i}$ be a function where each $f_{i}:2^{E}\to\mathbb{R}_{+}$ is a monotone submodular function. Suppose $\mathcal{P}\subseteq[0,1]^{n}$ is a polytope and define $\texttt{OPT}=\max_{\mathbf{x}\in\mathcal{P}}\widehat{F}(\mathbf{x})$. Then for any $\mathbf{x}\in[0,1]^{n}$ and $\mathbf{v}_{i}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{f}_{i}(\mathbf{x})\rangle$ we have
$\texttt{OPT}-\widehat{F}(\mathbf{x})\leq\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i},\nabla\widehat{f_{i}}(\mathbf{x})\rangle.$
Proof of Lemma A.1.
First, let us derive some observations on the decomposable submodular function $F$ and its multilinear extension $\widehat{F}$. By definition $\widehat{F}(\mathbf{x})=\mathbb{E}_{X\sim\mathbf{x}}[F(X)]$. Then
$\widehat{F}(\mathbf{x})=\mathbb{E}_{X\sim\mathbf{x}}[F(X)]=\mathbb{E}_{X\sim\mathbf{x}}\left[\sum_{i=1}^{N}p_{i}f_{i}\right]=\sum_{i=1}^{N}p_{i}(\mathbb{E}_{X\sim\mathbf{x}}[f_{i}(X)])=\sum_{i=1}^{N}p_{i}\widehat{f_{i}}(\mathbf{x})$
Then the gradient of $\widehat{F}$ at any point is equal to the average of gradients from local functions. That is $\nabla\widehat{F}(\mathbf{x})=\nabla\left(\sum_{i=1}^{N}p_{i}\widehat{f_{i}}(\mathbf{x})\right)=\sum_{i=1}^{N}p_{i}\nabla\widehat{f_{i}}(\mathbf{x}).$
Now let $\mathbf{w}\in\mathcal{P}$ be such that $\widehat{F}(\mathbf{w})=\texttt{OPT}$.
$\langle\mathbf{w},\nabla\widehat{F}(\mathbf{x})\rangle\leq\max_{\mathbf{v}\in\mathcal{P}}\langle\mathbf{v},\nabla\widehat{F}(\mathbf{x})\rangle=\max_{\mathbf{v}\in\mathcal{P}}\langle\mathbf{v},\sum_{i=1}^{N}p_{i}\nabla\widehat{f}_{i}(\mathbf{x})\rangle=\max_{\mathbf{v}\in\mathcal{P}}\sum_{i=1}^{N}p_{i}\langle\mathbf{v},\nabla\widehat{f_{i}}(\mathbf{x})\rangle$
$\leq\sum_{i=1}^{N}p_{i}\max_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{f_{i}}(\mathbf{x})\rangle=\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i},\nabla\widehat{f_{i}}(\mathbf{x})\rangle$
In what follows, we prove $\texttt{OPT}-\widehat{F}(\mathbf{x})\leq\langle\mathbf{w},\nabla\widehat{F}(\mathbf{x})\rangle$. Define $\mathbf{d}=(\mathbf{x}\vee\mathbf{w})-\mathbf{x}=(\mathbf{w}-\mathbf{x})\vee\mathbf{0}$. By the monotonicity, we have
$\texttt{OPT}=\widehat{F}(\mathbf{w})\leq\widehat{F}(\mathbf{x}\vee\mathbf{w}).$
Note that $\mathbf{d}>\mathbf{0}$, and hence by concavity of $\widehat{F}$ along any positive direction we get
$\widehat{F}(\mathbf{x}\vee\mathbf{w})=\widehat{F}(\mathbf{x}+\mathbf{d})\leq\widehat{F}(\mathbf{x})+\langle\mathbf{d},\nabla\widehat{F}(\mathbf{x})\rangle.$
Combining the above inequalities we obtain
$\texttt{OPT}-\widehat{F}(\mathbf{x})\leq\widehat{F}(\mathbf{x}\vee\mathbf{w})-\widehat{F}(\mathbf{x})\leq\langle\mathbf{d},\nabla\widehat{F}(\mathbf{x})\rangle$
Now, since $\nabla\widehat{F}(\mathbf{x})$ is nonnegative and $\mathbf{d}\leq\mathbf{w}$ then we get
$\texttt{OPT}-\widehat{F}(\mathbf{x})\leq\langle\mathbf{w},\nabla\widehat{F}(\mathbf{x})\rangle$
This completes the proof. ∎
The following is an immediate consequence of Lemma A.1.
Corollary A.2.
For any polytope $\mathcal{P}\subseteq[0,1]^{n}$, $\mathbf{x}^{(t)}\in[0,1]^{n}$, and $\mathbf{v}_{i}^{(t)}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle$, $\texttt{OPT}-\widehat{F}(\mathbf{x}^{(t)})\leq\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle.$
Next we use Assumption 3.1 to bound the difference between $\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle$ and $\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ at every iteration $t$.
Lemma A.3 (Bounded heterogeneity).
Let $\mathcal{M}$ be a matroid of rank $r$ and $\mathcal{P}$ be its corresponding matroid polytope. For any $\mathbf{x}^{(t)}\in[0,1]^{n}$ let $\mathbf{v}_{i}^{(t)}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$ and $\bar{\mathbf{v}}^{(t)}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)}$. Then, under the bounded gradient dissimilarity assumption 3.1, we have
$\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\geq\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-r\gamma_{t}$
Proof of Lemma A.3.
The proof is straightforward.
$\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$
$=\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\leq\sum_{i=1}^{N}p_{i}r\gamma_{t}=r\gamma_{t}$
where in the last inequality we used the fact that each $\mathbf{v}_{i}^{(t)}$ corresponds to a base in the matroid and we have $\mathbf{v}_{i}^{(t)}\in\{0,1\}^{n}$ and $|\mathbf{v}_{i}^{(t)}|_{1}=r$. Given this and Assumption 3.1 we have $|\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|\leq r\gamma_{t}$. ∎
A.1 Putting everything together, proof of Theorem 3.2.
We now are ready to prove Theorem 3.2.
Proof of Theorem 3.2.
Recall that $\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\Delta^{(t)}$ where in the full participation case $\Delta^{(t)}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)}$. According to the Taylor’s expansion of the function $\widehat{F}$ near the point $\mathbf{x}^{(t)}$ we can write
$\widehat{F}(\mathbf{x}^{(t+1)})$ $=\widehat{F}(\mathbf{x}^{(t)})+\langle\mathbf{x}^{(t+1)}-\mathbf{x}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle+\frac{1}{2}\langle\mathbf{x}^{(t+1)}-\mathbf{x}^{(t)},\mathbf{H}_{\widehat{F}}(\mathbf{x}^{(t+1)}-\mathbf{x}^{(t)})\rangle$ (12)
$=\widehat{F}(\mathbf{x}^{(t)})+\eta\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle+\frac{\eta^{2}}{2}\langle\Delta^{(t)},\mathbf{H}_{\widehat{F}}\Delta^{(t)}\rangle$ (13)
where $\mathbf{H}_{\widehat{F}}$ is the Hessian matrix i.e., the second derivative matrix. We provide a lower bound on each entry of $\mathbf{H}_{\widehat{F}}$. By the result of [Călinescu et al., 2011] and definition of the multilinear extension:
$\frac{\partial\widehat{F}}{\partial\mathbf{x}(i)\partial\mathbf{x}(j)}$ $=\mathbb{E}_{R\sim\mathbf{x}}[F(R\cup\{i,j\})-F(R\cup\{i\}\setminus\{j\})]-\mathbb{E}_{R\sim\mathbf{x}}[F(R\cup\{j\}\setminus\{i\})-F(R\setminus\{i,j\})]$ (14)
$\geq-\max\{F(\{i\}),F(\{j\})\}\geq-\max_{e\in E}F(\{e\})=-m_{F}$ (15)
where the second last inequality is a direct consequence of the submodularity of $F$. This means every entry of the Hessian is at least $-m_{F}$. Thus, we arrive at the following lower bound
$\langle\Delta^{(t)},\mathbf{H}_{\widehat{F}}\Delta^{(t)}\rangle\geq\sum_{i=1}^{n}\sum_{j=1}^{n}\Delta^{(t)}(i)\Delta^{(t)}(j)\mathbf{H}_{\widehat{F}}{(i,j)}\geq-m_{F}\sum_{i=1}^{n}\sum_{j=1}^{n}\Delta^{(t)}(i)\Delta^{(t)}(j)=-m_{F}\left(\sum_{i=1}^{n}\Delta^{(t)}(i)\right)^{2}\geq-m_{F}r^{2}$ (16)
where in the last inequality we used the fact that $\Delta^{(t)}\in\mathcal{P}$ as it is a convex combination of vectors from $\mathcal{P}$, and hence $\sum_{i=1}^{n}\Delta^{(t)}(i)\leq r$.
Thus from (13) and (16) it follows that
$\widehat{F}(\mathbf{x}^{(t+1)})$ $\geq\widehat{F}(\mathbf{x}^{(t)})+\eta\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (17)
It is now required to provide a lower bound for $\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)}\rangle$ in terms of OPT. We prove the following lemma.
Lemma A.4.
Let $\gamma_{t}$ be as in Assumption 3.1. Then we have $\texttt{OPT}-\widehat{F}(\mathbf{x}^{(t)})\leq\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)}\rangle+r\gamma_{t}$.
Proof of Lemma A.4.
By Corollary A.2 we have that
$\texttt{OPT}-\widehat{F}(\mathbf{x}^{(t)})\leq\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle$
Furthermore, by Lemma A.3, we have
$\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\leq r\gamma_{t}$
These two give the desired result. Moreover, it is worth noting
$\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\geq 0$
This is because
$\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\sum_{i=1}^{N}p_{i}\langle\Delta^{(t)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$
$=\sum_{i=1}^{N}p_{i}\left(\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\Delta^{(t)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right)\geq 0$
where in the last inequality by definition of $\mathbf{v}_{i}^{(t)}$ we have $\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\Delta^{(t)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\geq 0$. ∎
Followed by Lemma A.4 and equation 17, we obtain
$\widehat{F}(\mathbf{x}^{(t+1)})-\widehat{F}(\mathbf{x}^{(t)})$ $\geq\eta\left(\texttt{OPT}-\widehat{F}(\mathbf{x}^{(t)})-r\gamma_{t}\right)-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (18)
Now, by changing signs and adding OPT to both sides, we get
$\texttt{OPT}-\widehat{F}(\mathbf{x}^{(t+1)})$ $\leq(1-\eta)\left(\texttt{OPT}-\widehat{F}(\mathbf{x}^{(t)})\right)+\eta r\gamma_{t}+\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (19)
Applying the same inequality inductively gives
${\texttt{OPT}}-\widehat{F}(\mathbf{x}^{(t+1)})$ $\leq(1-\eta)^{t+1}\left({\texttt{OPT}}-\widehat{F}(\mathbf{0})\right)+\eta r\left(\sum_{t=0}^{T-1}(1-\eta)^{T-t-1}\gamma_{t}\right)+\frac{(\eta^{2}\cdot m_{F}r^{2})(\sum_{t=0}^{T-1}(1-\eta)^{T-t-1})}{2}$ (20)
$\leq(1-\eta)^{t+1}\left({\texttt{OPT}}-\widehat{F}(\mathbf{0})\right)+\eta r\left(\sum_{t=0}^{T-1}\gamma_{t}\right)+\frac{T\eta^{2}\cdot m_{F}r^{2}}{2}$ ($(1-\eta)<1$)
Hence for $\mathbf{x}^{(T)}$ we have
$\left(1-(1-\eta)^{T}\right){\texttt{OPT}}\leq\widehat{F}(\mathbf{x}^{(T)})+\eta r\left(\sum_{t=0}^{T-1}\gamma_{t}\right)+\frac{T\eta^{2}\cdot m_{F}r^{2}}{2}$ (21)
Finally, setting $\eta=1/T$ yields
$\left(1-\frac{1}{\mathrm{e}}\right)\texttt{OPT}$ $\leq\widehat{F}(\mathbf{x}^{(T)})+\frac{r\left(\sum_{t=0}^{T-1}\gamma_{t}\right)}{T}+\frac{m_{F}r^{2}}{2T}$ (22)
$=\widehat{F}(\mathbf{x}^{(T)})+\frac{rD}{T}+\frac{m_{F}r^{2}}{2T}$ (23)
for $\sum_{t=0}^{T-1}\gamma_{t}=D$. Note that for any $\varepsilon\geq 0$ by setting $T=\max\{\frac{2rD}{\varepsilon},\frac{m_{F}r^{2}}{\varepsilon}\}$ we obtain
$\left(1-\frac{1}{\mathrm{e}}\right)\texttt{OPT}\leq\widehat{F}(\mathbf{x}^{(T)})+\varepsilon$ (24)
Appendix B Proof of convergence for FedCG (Theorem 3.5)
In order to prove the approximation guarantee of Theorem 3.5, several steps are taken. First we show that our client sampling scheme is unbiased and furthermore we provide an upper bound for the variance.
B.1 Bounding the variance
Unbiased client selection.
Here we prove that our client selection is unbiased (Lemma 3.3).
Proof of Lemma 3.3.
Recall that $\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\langle\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ where $A^{(t)}$ is a set of size $K$.
$\mathbb{E}_{A^{(t)}}\left[\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\mathbb{E}_{A^{(t)}}\left[\langle\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]$ (25)
$=\frac{1}{K}\mathbb{E}_{A^{(t)}}\left[\langle\sum_{i\in A^{(t)}}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\frac{1}{K}\left[\sum_{i\in A^{(t)}}\mathbb{E}_{A^{(t)}}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]$ (26)
$=\frac{1}{K}\left[K\mathbb{E}_{A^{(t)}}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]$ (for an arbitrary $i\in A^{(t)}$)
$=\mathbb{E}_{A^{(t)}}\left[\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\langle\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (27)
Bounding the variance.
Bounding the variance is at the core of our proof of convergence. Recall the definition of $\gamma_{t}$ in Assumption 3.1 which plays a pivotal role in bounding the variance.
Proof of Lemma 3.4.
Recall $\Delta^{(t)}=\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\mathbf{v}_{i}^{(t)}$ and $|A^{(t)}|=K$ and let $\bar{\mathbf{v}}^{(t)}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)}$. By Lemma 3.3 we have that $\mathbb{E}_{A^{(t)}}\left[\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\langle\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$. Furthermore, our client sampling samples $K$ clients independently and with replacement. Therefore,
$\mathrm{Var}\left(\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right)$ $=\frac{1}{K}\mathbb{E}_{A^{(t)}}\left[\left(\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right)^{2}\right]$ (28)
where $\mathbf{s}$ corresponds to an arbitrary client in $A^{(t)}$. Note that both terms $\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ and $\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ are nonnegative. We provide an upper bound on the absolute value of $|\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|$ by considering two cases.
Case 1. In the first case we have $\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\geq\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$. Therefore,
$|\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|$ (29)
$=\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\sum_{i=1}^{N}p_{i}\langle\mathbf{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-\langle\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (30)
$\leq\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-\langle\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (31)
$\leq\sum_{i=1}^{N}p_{i}r\gamma_{t}=r\gamma_{t}$ (32)
where in the last inequality we used the fact that for each $\mathbf{v}_{i}^{(t)}$ we have $\mathbf{v}_{i}^{(t)}\in\{0,1\}^{n}$ and $|\mathbf{v}_{i}^{(t)}|_{1}=r$, and Assumption 3.1 that together yield $|\langle\mathbf{v}_{i},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|\leq r\gamma_{t}$.
Case 2. In the first case we have $\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\leq\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$. Therefore,
$|\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|=\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (33)
$=\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle+\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle$ (34)
$=\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle+\left(\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right)$ (35)
$\leq\sum_{i=1}^{N}p_{i}\left|\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle\right|+\sum_{i=1}^{N}p_{i}\left(\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right)$ (36)
$\leq r\gamma+\sum_{i=1}^{N}p_{i}\left(\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right)$ (37)
where in the last inequality we used the same argument as the Case 1. Now it is left to provide an upper bound for $\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x})\rangle-\langle\mathbf{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$, which by definition is nonnegative.
$\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$ (38)
$=\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t))}\rangle-\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle+\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$ (39)
$=\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle+\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$ (40)
$\leq\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle+\langle\mathbf{s},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$ (41)
$=\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle+\langle\mathbf{s},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$ (42)
By Assumption 3.1 we know that $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$ and $|\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$. Therefore, $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})|_{\infty}\leq 2\gamma_{t}$. Knowing that both $\mathbf{v}_{i}^{(t)},\mathbf{s}\in\{0,1\}^{n}$ and $|\mathbf{v}_{i}^{(t)}|_{1}=|\mathbf{s}|_{1}=r$ yields
$\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$ (43)
$\leq\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle+\langle\mathbf{s},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\leq 4r\gamma_{t}$ (44)
Putting together the above inequality and the inequality in equation 37 we obtain the following upper bound for Case 2
$|\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|\leq 5r\gamma_{t}.$
Provided the upper bounds in both cases we have
$\mathrm{Var}\left(\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right)$ $=\frac{1}{K}\mathbb{E}_{A^{(t)}}\left[\left(\langle\mathbf{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\bar{\mathbf{v}}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right)^{2}\right]\leq\frac{36r^{2}\gamma_{t}^{2}}{K}.$ (45)
B.2 Putting everything together, proof of Theorem 3.5.
We now are ready to prove Theorem 3.5.
Proof of Theorem 3.5.
Recall that $\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\Delta^{(t)}$. Similar to the proof of Theorem 3.2 equation (17) we derive that
$\widehat{F}(\mathbf{x}^{(t+1)})$ $\geq\widehat{F}(\mathbf{x}^{(t)})+\eta\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (46)
Given Lemmas 3.3 and 3.4 and using Chebyshev’s inequality, over the random choices of $A^{(t)}$ and for every $\alpha>0$ we obtain
$\mathbb{P}\left[\left|\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right|\leq\frac{6r\gamma_{t}/\sqrt{K}}{\alpha}\right]$ (47)
$\geq\mathbb{P}\left[\left|\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\mathbb{E}_{A^{(t)}}\left[\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]\right|\leq\frac{\sqrt{\mathrm{Var}\left(\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right)}}{\alpha}\right]\geq 1-\alpha^{2}$ (48)
Given this, with probability at least $1-\alpha^{2}$ and in the worst case it holds that
$\widehat{F}(\mathbf{x}^{(t+1)})$ $\geq\widehat{F}(\mathbf{x}^{(t)})+\eta\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (49)
$\geq\widehat{F}(\mathbf{x}^{(t)})+\eta\left(\langle\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\frac{6r\gamma_{t}/\sqrt{K}}{\alpha}\right)-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (50)
$\geq\widehat{F}(\mathbf{x}^{(t)})+\eta\left(\sum_{i=1}^{N}p_{i}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-r\gamma_{t}-\frac{6r\gamma_{t}/\sqrt{K}}{\alpha}\right)-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (by Lemma A.3)
$\geq\widehat{F}(\mathbf{x}^{(t)})+\eta\left(\texttt{OPT}-\widehat{F}(\mathbf{x}^{(t)})\right)-\eta\left(r\gamma_{t}+\frac{6r\gamma_{t}/\sqrt{K}}{\alpha}\right)-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (by Corollary A.2)
Now, by changing signs and adding OPT to both sides, we get
${\texttt{OPT}}-\widehat{F}(\mathbf{x}^{(t+1)})$ $\leq(1-\eta)\left({\texttt{OPT}}-\widehat{F}(\mathbf{x}^{(t)})\right)+\eta\left(r\gamma_{t}+\frac{6r\gamma_{t}/\sqrt{K}}{\alpha}\right)+\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (51)
Applying the same inequality inductively gives
${\texttt{OPT}}-\widehat{F}(\mathbf{x}^{(t+1)})$ $\leq(1-\eta)^{t+1}\left({\texttt{OPT}}-\widehat{F}(\mathbf{0})\right)+\eta\left(r\sum_{t=0}^{T-1}(1-\eta)^{T-t-1}\gamma_{t}+\frac{6r\sum_{t=0}^{T-1}(1-\eta)^{T-t-1}\gamma_{t}}{\alpha\sqrt{K}}\right)$ (52)
$+\frac{\sum_{t=0}^{T-1}(1-\eta)^{T-t-1}\eta^{2}\cdot m_{F}r^{2}}{2}$ (53)
$\leq(1-\eta)^{t+1}{\texttt{OPT}}+\eta\left(r\sum_{t=0}^{T-1}\gamma_{t}+\frac{6r\sum_{t=0}^{T-1}\gamma_{t}}{\alpha\sqrt{K}}\right)+\frac{T\eta^{2}\cdot m_{F}r^{2}}{2}$ (54)
Taking the union bound over $T$ steps and $\alpha=\sqrt{\frac{\delta}{T}}$, with probability at least $1-T\cdot\alpha^{2}=1-\delta$, we get
$\left(1-(1-\eta)^{T}\right){\texttt{OPT}}\leq\widehat{F}(\mathbf{x}^{(T)})+\eta\left(r\sum_{t=0}^{T-1}\gamma_{t}+\frac{6r\sum_{t=0}^{T-1}\gamma_{t}}{\sqrt{K\delta/T}}\right)+\frac{T\eta^{2}\cdot m_{F}r^{2}}{2}$ (55)
Setting $\eta=1/T$ yields
$\left(1-1/\mathrm{e}\right)\texttt{OPT}$ $\leq\widehat{F}(\mathbf{x}^{(T)})+\frac{1}{T}\left(r\sum_{t=0}^{T-1}\gamma_{t}+\frac{6r\sum_{t=0}^{T-1}\gamma_{t}}{\sqrt{K\delta/T}}\right)+\frac{m_{F}r^{2}}{2T}$ (56)
$=\widehat{F}(\mathbf{x}^{(T)})+\frac{rD}{T}+\frac{6rD}{\sqrt{KT\delta}}+\frac{m_{F}r^{2}}{2T}$ (57)
for $\sum_{t=0}^{T-1}\gamma_{t}=D$. Note that for any $\varepsilon\geq 0$ by setting $T=\max\{\frac{3rD}{\varepsilon},\frac{3m_{F}r^{2}}{2\varepsilon},\frac{324r^{2}D^{2}}{K\delta\varepsilon}\}$ we obtain
$\left(1-\frac{1}{\mathrm{e}}\right)\texttt{OPT}\leq\widehat{F}(\mathbf{x}^{(T)})+\varepsilon$ (58)
Appendix C Proof of convergence for FedCG+ (Theorem 4.3)
C.1 Bounding the variance
Unbiased client selection.
Here we prove that our client selection is unbiased (Lemma 4.1). The proof is almost identical to the one for Lemma 3.3. We present a proof here for the sake of completeness.
Proof of Lemma 4.1.
Recall that $\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\langle\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ where $A^{(t)}$ is a set of size $K$.
$\mathbb{E}_{A^{(t)}}\left[\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\mathbb{E}_{A^{(t)}}\left[\langle\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]$ (59)
$=\frac{1}{K}\mathbb{E}_{A^{(t)}}\left[\langle\sum_{i\in A^{(t)}}\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\frac{1}{K}\left[\sum_{i\in A^{(t)}}\mathbb{E}_{A^{(t)}}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]$ (60)
$=\frac{1}{K}\left[K\mathbb{E}_{A^{(t)}}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]$ (for an arbitrary $i\in A^{(t)}$)
$=\mathbb{E}_{A^{(t)}}\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\langle\sum_{i=1}^{N}p_{i}\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (61)
Bounding the variance.
Bounding the variance is at the core of our proof of convergence. Assumption 3.1 plays a pivotal role in bounding the variance. The main difficulty and difference between this proof and the proof of Lemma 3.4 is that firstly the variance is effected by the divergence caused by local steps, and secondly $\widetilde{\Delta}_{i}^{(t+\tau)}$ may not be integral vectors and it could potentially have $O(n)$ nonzero entries.
In order to handle the divergence caused by local steps we assume Lipschitzness. First, let us derive useful inequalities using the Lipschitzness condition.
Consequences of Lipschitzness. Consider $t$-th iteration and recall the definition of $L_{t}$ i.e., $L_{t}$ is such that for all $i\in[N]$ and $j\in[\tau]$ it holds
$\left\|\nabla\widehat{f_{i}}(\mathbf{x}_{i}^{(t,j)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\right\|\leq L_{t}\|\mathbf{x}_{i}^{(t,j)}-\mathbf{x}^{(t)}\|$
where $\mathbf{x}_{i}^{(t)}=\mathbf{x}^{(t)}$ and $\mathbf{x}_{i}^{(t,j)}$ are the local models for client $i$ at time step $t$ and $t+j$, respectively. We first bound the divergence of $\mathbf{x}_{i}^{(t,\tau)}$ from $\mathbf{x}^{(t)}$.
$\left\|\mathbf{x}_{i}^{(t,\tau)}-\mathbf{x}^{(t)}\right\|=\left\|\frac{1}{\tau}\sum_{j=0}^{\tau-1}\widetilde{\mathbf{v}}_{i}^{(t,j)}\right\|\leq\frac{1}{\tau}\sum_{j=0}^{\tau-1}\left\|\widetilde{\mathbf{v}}_{i}^{(t,j)}\right\|\leq\sqrt{r}$ (62)
Since $\widetilde{\mathbf{v}}_{i}^{(t,j)}\in\mathcal{P}$, the same upper bound holds for every $0\leq j\leq\tau$; $\|\mathbf{x}_{i}^{(t,j)}-\mathbf{x}^{(t)}\|\leq\sqrt{r}$. Assuming $L_{t}$, for all $0\leq j\leq\tau$,
$\left\|\nabla\widehat{f_{i}}(\mathbf{x}_{i}^{(t,j)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\right\|\leq L\|\mathbf{x}_{i}^{(t,j)}-\mathbf{x}^{(t)}\|\leq L_{t}\sqrt{r}$ (63)
Note $\zeta_{i}^{(t,j)}$ are sampled according to $\mathbf{x}_{i}^{(t,j)}$, and $\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})$ estimates $\nabla\widehat{f_{i}}(\mathbf{x}_{i}^{(t,j)})$ within factor $\sigma$ i.e., $\|\nabla\widehat{f_{i}}(\mathbf{x}_{i}^{(t,j)})-\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})\|\leq\sigma$. Hence, given this estimation and equation (63), for every $0\leq j\leq\tau$ it holds that
$\left\|\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\right\|\leq\sigma+L_{t}\sqrt{r}$ (64)
Proof of Lemma 4.2.
By Lemma 4.1 we know:
$\mathbb{E}_{A^{(t)}}\left[\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\langle\sum_{i=1}^{N}p_{i}\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$
where $\widetilde{\Delta}^{(t+\tau)}=\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\widetilde{\Delta}_{i}^{(t+\tau)}$. Define $\overline{\Delta}^{(t+\tau)}=\sum_{i=1}^{N}p_{i}\widetilde{\Delta}_{i}^{(t+\tau)}$. Each client is selected to be in set $A^{(t)}$ independently and with replacement. Therefore,
$\mathrm{Var}(\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle)=\frac{1}{K}\mathbb{E}_{A^{(t)}}\left[\left(\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right)^{2}\right]$ (65)
where $\widetilde{\Delta}_{s}^{(t+\tau)}=\mathbf{x}_{s}^{(t,\tau)}-\mathbf{x}_{s}^{(t,0)}$ corresponds to an arbitrary client in $A^{(t)}$. Note that both terms $\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ and $\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ are nonnegative. We provide an upper bound on the absolute value of $\left|\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right|$ by considering two cases.
Case 1. In the first case we have $\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\geq\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$. Therefore,
$\left|\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right|$ (66)
$=\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (67)
$=\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-\langle\sum_{i=1}^{N}p_{i}\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (68)
Observe that $\widetilde{\Delta}_{s}^{(t+\tau)}=\frac{1}{\tau}\sum_{j=0}^{\tau-1}\widetilde{\mathbf{v}}_{s}^{(t,j)}$ with $\widetilde{\mathbf{v}}_{s}^{(t,j)}=\operatornamewithlimits{argmax}_{\mathbf{v}\in P(M)}\langle\mathbf{v},\nabla\widetilde{f}_{s}(\mathbf{x}_{s}^{(t,j)},\zeta_{s}^{(t,j)})\rangle$.
Therefore, using the Lipschitzness condition we get the following. In what follows let $\mathbf{d}_{1}=L_{t}\sqrt{r}\mathbf{1}$ and $\mathbf{d}_{2}=\sigma\mathbf{1}$ be vectors of length $n$ where every components are $L_{t}\sqrt{r}$ and $\sigma$, respectively.
$\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t)})\rangle=\frac{1}{\tau}\sum_{j=0}^{\tau-1}\langle\widetilde{\mathbf{v}}_{s}^{(t,j)},\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t)})\rangle$ (69)
$\leq\frac{1}{\tau}\sum_{j=0}^{\tau-1}\langle\widetilde{\mathbf{v}}_{s}^{(t,j)},\nabla\widetilde{f}_{i}(\mathbf{x}_{i}^{(t,j)})+\mathbf{d}_{1}+\mathbf{d}_{2}\rangle$ (by equation (64))
$\leq\frac{1}{\tau}\sum_{j=0}^{\tau-1}\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widetilde{f}_{i}(\mathbf{x}_{i}^{(t,j)})+\mathbf{d}_{1}+\mathbf{d}_{2}\rangle$ (by definition of $\widetilde{\mathbf{v}}_{i}^{(t,j)}$)
$\leq\frac{1}{\tau}\sum_{j=0}^{\tau-1}\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t)})\rangle+\frac{1}{\tau}\sum_{j=0}^{\tau-1}\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\mathbf{d}_{1}+\mathbf{d}_{2}\rangle$ (70)
$\leq\frac{1}{\tau}\sum_{j=0}^{\tau-1}\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t)})\rangle+(\sigma r+L_{t}r^{1.5})$ (71)
$=\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t)})\rangle+(\sigma r+L_{t}r^{1.5})$ (72)
$=\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle+(\sigma r+L_{t}r^{1.5})$ ($\mathbf{x}_{i}^{(t)}=\mathbf{x}^{(t)}$)
Therefore, for Case 1 we get
$\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-\langle\sum_{i=1}^{N}p_{i}\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (73)
$\leq\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-\langle\sum_{i=1}^{N}p_{i}\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle+(\sigma r+L_{t}r^{1.5})$ (74)
$=\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x})-\nabla\widehat{F}(\mathbf{x})\rangle+(\sigma r+L_{t}r^{1.5})$ (75)
$=\sum_{i=1}^{N}p_{i}\left(\frac{1}{\tau}\sum_{j=0}^{\tau-1}\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widehat{f_{i}}(\mathbf{x})-\nabla\widehat{F}(\mathbf{x})\rangle\right)+(\sigma r+L_{t}r^{1.5})$ (76)
$\leq\sum_{i=1}^{N}p_{i}\left(\frac{1}{\tau}\sum_{j=0}^{\tau-1}r\gamma_{t}\right)+(\sigma r+L_{t}r^{1.5})$ (Assumption 3.1)
$\leq r\gamma_{t}+(\sigma r+L_{t}r^{1.5})$ (77)
Note that in the above we used the fact that for each $\widetilde{\mathbf{v}}_{i}^{(t,j)}$ we have $\widetilde{\mathbf{v}}_{i}^{(t,j)}\in\{0,1\}^{n}$ and $|\widetilde{\mathbf{v}}_{i}^{(t,j)}|_{1}=r$.
Case 2. In this case we have $\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\leq\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$. Therefore,
$|\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|$ (78)
$=\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (79)
$=\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ (80)
$+\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$ (81)
$\leq\sum_{i=1}^{N}p_{i}\left|\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle\right|$ (82)
$+\sum_{i=1}^{N}p_{i}\left(\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right)$ (83)
$\leq\sum_{i=1}^{N}\frac{p_{i}}{\tau}\sum_{j=0}^{\tau-1}\left|\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widehat{F}(\mathbf{x}^{(t)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle\right|$ (84)
$+\sum_{i=1}^{N}p_{i}\left(\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right)$ (85)
$\leq r\gamma_{t}+\sum_{i=1}^{N}p_{i}\left(\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right)$ (86)
where in the last inequality we used the same argument by noting $\widetilde{\mathbf{v}}_{i}^{(t,j)}\in\{0,1\}^{n}$ and $|\widetilde{\mathbf{v}}_{i}^{(t,j)}|_{1}=r$. Now it is left to provide an upper bound for $\left|\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right|$.
$\left|\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right|$ (87)
$=\left|\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle+\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right|$ (88)
$=\left|\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle+\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right|$ (89)
$\leq\left|\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle\right|+\left|\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right|$ (90)
By Assumption 3.1 we know that $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$ and $|\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})|_{\infty}\leq\gamma_{t}$. Therefore, $|\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})|_{\infty}\leq 2\gamma_{t}$. This is used to upper bound the first term in (90). In order to bound the second term in (90) we appeal to equation (72) in Case 1. (Note that there are two cases two consider here because of the absolute value, however the argument is similar and we argue about one case.) Let $\mathbf{1}$ be the all one vector of length $n$, then
$\left|\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle\right|+\left|\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right|$ (91)
$\leq\frac{1}{\tau}\sum_{j=0}^{\tau-1}\left|\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle\right|+\left|\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right|$ (92)
$+(\sigma r+L_{t}r^{1.5})$ (93)
$\leq\frac{1}{\tau}\sum_{j=0}^{\tau-1}\left|\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})-\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})\rangle\right|+\frac{1}{\tau}\sum_{j=0}^{\tau-1}\left|\langle\widetilde{\mathbf{v}}_{s}^{(t,j)},\nabla\widehat{f_{s}}(\mathbf{x}^{(t)})-\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\right|$ (94)
$+(\sigma r+L_{t}r^{1.5})$ (95)
$\leq\frac{1}{\tau}\sum_{j=0}^{\tau-1}\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},2\gamma_{t}\mathbf{1}\rangle+\frac{1}{\tau}\sum_{j=0}^{\tau-1}\langle\widetilde{\mathbf{v}}_{s}^{(t,j)},2\gamma_{t}\mathbf{1}\rangle+(\sigma r+L_{t}r^{1.5})$ (96)
$\leq 4r\gamma_{t}+(\sigma r+L_{t}r^{1.5})$ (97)
where in the last inequality we used the fact that each $\widetilde{\mathbf{v}}_{i}^{(t,j)},\widetilde{\mathbf{v}}_{s}^{(t,j)}\in\{0,1\}^{n}$ and $|\widetilde{\mathbf{v}}_{i}^{(t,j)}|_{1}=|\widetilde{\mathbf{v}}_{s}^{(t,j)}|_{1}=r$.
Putting together the above inequality and the inequality in equation 86 we obtain the following upper bound for Case 2
$\left|\langle\widetilde{\Delta}_{s}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right|\leq 5r\gamma_{t}+(\sigma r+L_{t}r^{1.5}).$
Finally, we are at the place where we can present our upper bound for the variance
$\mathrm{Var}(\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle)$ $\leq\frac{1}{K}\mathbb{E}_{A^{(t)}}\left[\left(6r\gamma+2(\sigma r+Lr^{1.5})\right)^{2}\right]$
$=\frac{1}{K}\left(6r\gamma_{t}+2(\sigma r+L_{t}r^{1.5})\right)^{2}.\qed$
C.2 Proof of Theorem 4.3
Convergence of FedCG+.
While at the high level this proof is similar to the previous convergence proofs in Theorems 3.2, 3.5, it still requires taking care of the error caused due to the local steps. (This is an additive error term that cannot be controlled by sampling more clients at each round.)
Proof of Theorem 4.3.
For any $t\in\mathcal{I}_{\tau}$ we analyze the difference between $\widehat{F}(\mathbf{x}^{(t)})$ and $\widehat{F}(\mathbf{x}^{(t+\tau)})$. Recall that $\mathbf{x}^{(t+\tau)}\leftarrow\mathbf{x}^{(t)}+\eta\widetilde{\Delta}^{(t+\tau)}$ where $\eta$ is the server’s learning rate and $\widetilde{\Delta}^{(t+\tau)}=\frac{1}{|A^{(t)}|}\sum_{i\in A^{(t)}}\widetilde{\Delta}_{i}^{(t+\tau)}=\frac{1}{\tau|A^{(t)}|}\sum_{i\in A^{(t)}}\sum_{j=0}^{\tau-1}\widetilde{\mathbf{v}}_{i}^{(t,j)}$. Similar to the proof of Theorem 3.2 equation (17) we derive that
$\widehat{F}(\mathbf{x}^{(t+\tau)})-\widehat{F}(\mathbf{x}^{(t)})$ $\geq\eta\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)}\rangle-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (98)
Let $\overline{\Delta}^{(t+\tau)}=\sum_{i=1}^{N}p_{i}\widetilde{\Delta}_{i}^{(t+\tau)}$ and by Lemma 4.1 we have that $\mathbb{E}_{A^{(t)}}\left[\eta\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right]=\eta\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$. Given Lemma 4.2 and using Chebyshev’s inequality, over the random choices of $A^{(t)}$ and for every $\alpha>0$ and $\chi^{2}=\frac{1}{K}\left(6r\gamma_{t}+2(\sigma r+L_{t}r^{1.5})\right)^{2}$ we obtain
$\mathbb{P}\left[\left|\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\right|\leq\frac{\chi}{\alpha}\right]\geq 1-\alpha^{2}$ (99)
Given this and (98), with probability at least $1-\alpha^{2}$ and in the worst case it holds that
$\widehat{F}(\mathbf{x}^{(t+\tau)})$ $\geq\widehat{F}(\mathbf{x}^{(t)})+\eta\langle\widetilde{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)}\rangle-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (100)
$\geq\widehat{F}(\mathbf{x}^{(t)})+\eta\left(\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\frac{\chi}{\alpha}\right)-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (101)
Claim C.1.
Let $\gamma_{t}$ be as in Assumption 3.1, we have $\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\geq\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-r\gamma_{t}$.
Proof.
Proof is similar to the proof of Lemma A.3.
$\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-\langle\overline{\Delta}^{(t+\tau)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle=\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$
$=\sum_{i=1}^{N}\sum_{j=0}^{\tau-1}\frac{p_{i}}{\tau}\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})-\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle\leq\sum_{i=1}^{N}\sum_{j=0}^{\tau-1}\frac{p_{i}}{\tau}r\gamma_{t}=r\gamma_{t}\qed$
Therefore,
$\widehat{F}(\mathbf{x}^{(t+\tau)})$ $\geq\widehat{F}(\mathbf{x}^{(t)})+\eta(\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-r\gamma_{t}-\frac{\chi}{\alpha})-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (102)
Claim C.2.
$\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle\geq\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-2(\sigma r+L_{t}r^{1.5})$.
Hence, by Claim C.2 and equation (102) we have
$\widehat{F}(\mathbf{x}^{(t+\tau)})$ $\geq\widehat{F}(\mathbf{x}^{(t)})+\eta(\sum_{i=1}^{N}p_{i}\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-r\gamma_{t}-\frac{\chi}{\alpha}-2(\sigma r+L_{t}r^{1.5}))-\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (103)
Now, by changing signs and adding OPT to both sides, we get
${\texttt{OPT}}-\widehat{F}(\mathbf{x}^{(t+\tau)})$ $\leq(1-\eta)\left({\texttt{OPT}}-\widehat{F}(\mathbf{x}^{(t)})\right)+\eta\left(r\gamma_{t}+\frac{\chi}{\alpha}+2(\sigma r+L_{t}r^{1.5})\right)+\frac{\eta^{2}\cdot m_{F}r^{2}}{2}$ (104)
Applying the same inequality inductively gives and using $1>1-\eta$
${\texttt{OPT}}-\widehat{F}(\mathbf{x}^{(t+\tau)})$ $\leq(1-\eta)^{t+1}\left({\texttt{OPT}}-\widehat{F}(\mathbf{0})\right)$ (105)
$+\eta\Big{(}r\sum_{t\in I_{\tau}}\gamma_{t}+\frac{6r\sum_{t\in I_{\tau}}\gamma_{t}+2(\sigma rT/\tau+\sum_{t\in I_{\tau}}L_{t}r^{1.5})}{\alpha\sqrt{K}}+2(\sigma rT/\tau+r^{1.5}\sum_{t\in I_{\tau}}L_{t})\Big{)}+\frac{T\eta^{2}\cdot m_{F}r^{2}}{2}$ (106)
$=(1-\eta)^{t+1}{\texttt{OPT}}$ (107)
$+\eta\Big{(}r\sum_{t\in I_{\tau}}\gamma_{t}+\frac{6r\sum_{t\in I_{\tau}}\gamma_{t}+2(\sigma rT/\tau+\sum_{t\in I_{\tau}}L_{t}r^{1.5})}{\alpha\sqrt{K}}+2(\sigma rT/\tau+r^{1.5}\sum_{t\in I_{\tau}}L_{t})\Big{)}+\frac{T\eta^{2}\cdot m_{F}r^{2}}{2}$ (108)
Taking the union bound over $T/\tau$ steps and $\alpha=\sqrt{\frac{\delta\tau}{T}}$, with probability at least $1-T\cdot\alpha^{2}/\tau=1-\delta$, we get
$\left(1-(1-\eta)^{T/\tau}\right){\texttt{OPT}}$ $\leq\widehat{F}(\mathbf{x}^{(T/\tau)})+\eta\left(r\sum_{t\in I_{\tau}}\gamma_{t}+\frac{6r\sum_{t\in I_{\tau}}\gamma_{t}+2(\sigma rT/\tau+\sum_{t\in I_{\tau}}L_{t}r^{1.5})}{\sqrt{K\tau\delta/T}}+2(\sigma rT/\tau+r^{1.5}\sum_{t\in I_{\tau}}L_{t})\right)$ (109)
$+\frac{T\eta^{2}\cdot m_{F}r^{2}}{2\tau}$ (110)
Setting $\eta=\tau/T$ yields
$\left(1-1/\mathrm{e}\right)\texttt{OPT}$ $\leq\widehat{F}(\mathbf{x}^{(T/\tau)})+\frac{\tau}{T}\left(r\sum_{t\in I_{\tau}}\gamma_{t}+\frac{6r\sum_{t\in I_{\tau}}\gamma_{t}+2(\sigma rT/\tau+\sum_{t\in I_{\tau}}L_{t}r^{1.5})}{\sqrt{K\tau\delta/T}}+2(\sigma rT/\tau+r^{1.5}\sum_{t\in I_{\tau}}L_{t})\right)$ (111)
$+\frac{\tau\cdot m_{F}r^{2}}{2T}$ (112)
$=\widehat{F}(\mathbf{x}^{(T/\tau)})+\frac{\tau rD}{T}+\frac{\sqrt{\tau}(6rD+2r^{1.5}Q)}{\sqrt{KT\delta}}+2\sigma r(\frac{\sqrt{\tau}}{\sqrt{KT\delta}}+1)+\frac{2\tau r^{1.5}Q}{T}+\frac{\tau\cdot m_{F}r^{2}}{2T}$ (113)
Proof of Claim C.2
Proof of Claim C.2.
Recall the definitions of $\widetilde{\Delta}_{i}^{(t+\tau)}$ and $\mathbf{v}_{i}^{(t)}$. For $\widetilde{\Delta}_{i}^{(t+\tau)}=\frac{1}{\tau}\sum_{j=0}^{\tau-1}\widetilde{\mathbf{v}}_{i}^{(t,j)}$ where for each $j$ we have $\widetilde{\mathbf{v}}_{i}^{(t,j)}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widetilde{f}_{i}(\mathbf{x}_{i}^{(t,j)})\rangle$. Furthermore, $\mathbf{v}_{i}^{(t)}=\operatornamewithlimits{argmax}_{\mathbf{y}\in\mathcal{P}}\langle\mathbf{y},\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t)})\rangle$. We have,
$\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\Delta}_{i}^{(t+\tau)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle=\frac{1}{\tau}\sum_{j=0}^{\tau-1}(\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle-\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle)$ (114)
On one hand, both $\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t)})$ and $\nabla\widetilde{f}_{i}(\mathbf{x}_{i}^{(t,j)})$ are nonnegative vectors and each $\widetilde{\mathbf{v}}_{i}^{(t,j)}$ and $\mathbf{v}_{i}^{(t)}$ corresponds to a maximum-weight independent set in the matroid, with respect to the gradient vectors, and they can be found easily by a greedy algorithm. On the other hand, equation (64) tells us $\left\|\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})-\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\right\|\leq\sigma+L_{t}\sqrt{r}$.
Let $A=\{e\mid\mathbf{v}_{i}^{(t)}(e)=1\}$ be the set of indices where $\mathbf{v}_{i}^{(t)}$ is one, similarly define $B=\{e\mid\widetilde{\mathbf{v}}_{i}^{(t,j)}(e)=1\}$. Then, by definition and equation (64) we get:
$\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$ $=\sum_{a\in A}\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})(a)$ (115)
$\geq\sum_{b\in B}\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})(b)\geq\sum_{b\in B}\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})(b)-r(\sigma+L_{t}\sqrt{r})$ (116)
Similarly, we have
$\langle\widetilde{\mathbf{v}}_{i}^{(t,j)},\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})\rangle$ $=\sum_{b\in B}\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})(b)$ (117)
$\geq\sum_{a\in A}\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})(a)\geq\sum_{a\in A}\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})(a)-r(\sigma+L_{t}\sqrt{r}).$ (118)
Putting the above two together gives us the desired bound. ∎
C.3 Gradient estimation
In terms of computation cost on clients’ devices, we point out that the definition of multilinear extension involves summing over all subsets $S$ of $E$. There are $2^{|E|}$ such subsets, thus even computing $\widehat{g}(\mathbf{x})$ for a single $\mathbf{x}$ could take exponential time. However, we can randomly sample $m$ subsets $R_{1},\dots,R_{m}$ of $E$ according to $\mathbf{x}$. Then a simple application of Chernoff’s bound shows for any multilinear extension $\widehat{g}$ and $\mathbf{x}$, $\left|\frac{1}{m}\sum_{i=1}^{m}g(R_{i})-\widehat{g}(\mathbf{x})\right|\leq\sigma\max_{S}g(S)$ [Vondrák, 2008, Călinescu et al., 2011] with probability at least $1-\mathrm{e}^{-m\sigma^{2}/4}$. Observe that $\frac{\partial\widehat{g}}{\partial\mathbf{x}(e)}=\mathbb{E}[g(R\cup\{e\})]-\mathbb{E}[g(R)]$ where $R$ is a random subset of $E\setminus\{e\}$ sampled according to $\mathbf{x}$. Hence, by a similar argument, with $m$ random samples, we can compute an $\sigma$-approximation of $\nabla\widehat{g}(\mathbf{x})$ with $1-\mathrm{e}^{-m\sigma^{2}/4}$ probability. By suitably choosing $m$ as in Algorithm 2, we can assume the gradients are estimated within $(1\pm\sigma)$ accuracy with high probability.
The above discussion yields:
Lemma C.3.
Let $\sigma>0$ be an error for gradient estimation and set $m=O(\log{(1/\delta)}/\sigma^{2})$ for $\delta>0$. Let $\zeta_{i}^{(t,j)}=\{R_{i}^{(t,j,1)},\dots,R_{i}^{(t,j,m)}\}$ be subsets of $E$ that are sampled independently according to $i$-th client’s local model $\mathbf{x}_{i}^{(t,j)}$. Let $\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})$ denote the stochastic gradient for the $i$-th client that approximates $\nabla\widehat{f_{i}}(\mathbf{x}_{i}^{(t,j)})$. Then with $1-\delta$ probability we have $\|\nabla\widehat{f_{i}}(\mathbf{x}_{i}^{(t,j)})-\nabla\widetilde{f_{i}}(\mathbf{x}_{i}^{(t,j)},\zeta_{i}^{(t,j)})\|\leq\sigma$.
Appendix D Discrete examples
We consider two well-studied problems, namely Facility Location and Maximum Coverage problems and discuss the details of how to compute importance factors in federated setting efficiently and prove that the expected number of clients participating in each round is small for these two problems.
D.0.1 Facility Location Problem
Let $\mathcal{C}$ be a set of $N$ clients and $E$ be a set of facilities with $|E|=n$. For $c:\mathcal{C}\times E\to\mathbb{R}$ let the $i$-th client’s score function over a subset of facilities be $f_{i}(A)=\max_{j\in A}c(i,j)$. The objective for Max Facility Location is
$\max_{S\subseteq E,|S|\leq k}\left\{F(A)=\sum_{i=1}^{N}\max_{j\in A}c(i,j)\right\}$
For each client $i$ the importance factor is $\max\limits_{j\in E}\frac{c(i,j)}{F(\{j\})}$.
Computing importance factors in federated setting.
Here we discuss that each client can compute its corresponding importance factor in a federated setting and using a secure aggregator without sharing its data with other clients. Because of the simple nature of importance factors for this problem, one can easily verify that in Algorithm 4 every client correctly computes its own importance factor. Furthermore, Algorithm 4 requires only two communication rounds.
Theorem D.1.
In Algorithm 4, every clients correctly computes its own importance factor. Moreover, Algorithm 4 has only two communication rounds and during each round each client requires only $O(n)$ local function evaluations.
Having $w_{i}$ on hand we can execute Algorithm 3 for Max Facility Location problem. Note that, in this problem we are dealing with a uniform matroid of rank $k$.
Theorem D.2.
Suppose clients’ importance factors are computed using Algorithm 4 and let $\varepsilon\in(0,1)$. Algorithm 3 after $k$ communication rounds returns a set $S$ of size $k$ such that with probability at least $1-1/n$
$(1-1/\mathrm{e}-\varepsilon)\texttt{OPT}\leq F(S)$
Moreover, the expected number of clients participating during each round is $\tilde{O}(kn^{2}/\varepsilon^{2})$.
Proof.
The approximation guarantee follows from Theorem 5.1. The expected number of clients participating in each round of Algorithm 3 is $\tilde{O}(kn^{2}/\varepsilon^{2})$. This is because
$\sum_{i=1}^{n}\kappa_{i}\leq\kappa\sum_{i=1}^{n}w_{i}\leq\tilde{O}(kn/\varepsilon^{2})\sum_{i=1}^{N}w_{i}=\tilde{O}(kn/\varepsilon^{2})\sum_{i=1}^{N}\max\limits_{j\in E}\frac{c(i,j)}{F(\{j\})}\leq\tilde{O}(kn/\varepsilon^{2})\sum_{j=1}^{|E|}\frac{\sum_{i\in I}c(i,j)}{F(\{j\})}$
$=\tilde{O}(kn/\varepsilon^{2})\sum_{j=1}^{|E|}\frac{F(\{j\})}{F(\{j\})}=\tilde{O}(kn^{2}/\varepsilon^{2})$
Algorithm 4 Computing importance factors for Facility Location
1:  Input: Ground set $E$
2:  Let $\mathcal{O}$ be a vector of length $n$. {intention: $\mathcal{O}[i]=F(\{i\})$.}
3:  for each client $i$ in parallel do
4:     Compute $\mathcal{O}_{i}=[f_{i}(\{1\}),\dots,f_{i}(\{|E|\})]$
5:     Send $\mathcal{O}_{i}$ back to the secure aggregator.
6:  end for
7:  SecAgg: $\mathcal{O}=\sum_{i\in[N]}\mathcal{O}_{i}$
8:  Server sends $\mathcal{O}$ to all clients.
9:  for each client $i$ in parallel do
10:     Compute $w_{i}=\max_{j\in E}\frac{c(i,j)}{\mathcal{O}[j]}$
11:  end for
D.0.2 Maximum Coverage Problem
Let $\mathcal{C}=\{C_{1},\dots,C_{N}\}$ be a set of clients and $E=\{G_{1},\dots,G_{n}\}$ be a family of sets where each $G_{i}\subseteq\mathcal{C}$ is a group of clients. Given a positive integer $k$, in the Max Coverage problem the objective is to select at most $k$ groups of clients from $E$ such that the maximum number of clients are covered, i.e., the union of the selected groups has maximal size. One can formulate this problem as follows. For every $i\in[N]$ and $A\subseteq[n]$ define $f_{i}(A)$ as
$f_{i}(A)=\begin{cases}1&\text{if there exists $a\in A$ such that$C_{i}\in G_{a}$},\\ 0&\text{otherwise.}\end{cases}$
Note that $f_{i}$’s are monotone and submodular. Furthermore, define $F(A)=\sum_{i\in[N]}f_{i}(A)$ which is monotone and submodular as well. Now the Max Coverage problem is equivalent to
$\max_{A\subseteq[n],|A|\leq k}\left\{F(A)=\sum_{i\in[N]}f_{i}(A)\right\}$ (119)
For each client $C_{i}$, its importance factor $w_{i}$ is $\max\limits_{\begin{subarray}{c}G_{a}\in E,C_{i}\in G_{a}\end{subarray}}\frac{1}{|G_{a}|}$.
Computing importance factors in federated setting.
Here we discuss that each client can compute its corresponding importance factor in a federated setting and using a secure aggregator without sharing which group she belongs to. Because of the simple nature of importance factors for this problem, one can easily verify that in Algorithm 5 every client correctly computes its own importance factor. Furthermore, Algorithm 5 requires only two communication rounds.
Theorem D.3.
In Algorithm 5, every clients correctly computes its own importance factor without revealing which groups they belong to. Moreover, Algorithm 5 has only two communication rounds.
Remark D.4.
We point out that a simple algorithm where at each round clients share their membership with the server using SecAgg protocols is applicable here. However, such algorithm requires full client participation at each round. This can be resolved by sampling sufficiently many clients at each round. However, the number of sampled clients as well as the approximation factor would depend on $N$.
Having $w_{i}$ on hand we can execute Algorithm 3 for Max Coverage problem. Note that, in this problem we are dealing with a uniform matroid of rank $k$.
Theorem D.5.
Suppose clients’ importance factors are computed using Algorithm 5 and let $\varepsilon\in(0,1)$. Algorithm 3 after $k$ communication rounds returns a set $S\subseteq E$ of size $k$ such that with probability at least $1-1/n$
$(1-1/\mathrm{e}-\varepsilon)\texttt{OPT}\leq F(S)$
Moreover, the expected number of clients participating during each round is $\tilde{O}(kn^{2}/\varepsilon^{2})$.
Proof.
The approximation guarantee follows from Theorem 5.1. The expected number of clients participating in each round of Algorithm 3 is $\tilde{O}(kn^{2}/\varepsilon^{2})$. This is because
$\sum_{i=1}^{n}\kappa_{i}\leq\kappa\sum_{i=1}^{n}w_{i}\leq\tilde{O}(kn/\varepsilon^{2})\sum_{i=1}^{N}w_{i}=\tilde{O}(kn/\varepsilon^{2})\sum_{i=1}^{N}\max\limits_{\begin{subarray}{c}G_{a}\in E,C_{i}\in G_{a}\end{subarray}}\frac{1}{|G_{a}|}\leq\tilde{O}(kn/\varepsilon^{2})\sum_{j=1}^{|E|}\frac{|G_{j}|}{|G_{j}|}=\tilde{O}(kn^{2}/\varepsilon^{2})$
Algorithm 5 Computing importance factors for Max Coverage
1:  Let $\mathcal{O}$ be a vector of length $n$. {intention: $\mathcal{O}[i]=|G_{i}|$.}
2:  for each client $i$ in parallel do
3:     Compute vector $\mathcal{O}_{i}\in\{0,1\}^{n}$
$\begin{cases}\mathcal{O}_{i}[a]=1&\text{if }C_{i}\in G_{a}\\ \mathcal{O}_{i}[a]=0&\text{otherwise}\end{cases}$
4:     Send $\mathcal{O}_{i}$ back to the secure aggregator.
5:  end for
6:  Secure aggregator computes $\mathcal{O}=\sum_{i\in[N]}\mathcal{O}_{i}$ and sends it to the server.
7:  Server sends $\mathcal{O}$ to all clients.
8:  for each client $i$ in parallel do
9:     Compute $w_{i}=\max\limits_{\begin{subarray}{c}C_{i}\in G_{a}\end{subarray}}\frac{1}{|\mathcal{O}[a]|}$
10:  end for